### v1

# Improved Constructions of Reed-Solomon Codes with Optimal Repair Bandwidth

Abstract
Maximum-distance-separable (MDS) codes are widely used in distributed storage, yet naive repair of a single erasure in an $[n,k]$ MDS code downloads the entire contents of $k$ nodes. Minimum Storage Regenerating (MSR) codes (Dimakis et al., 2010) minimize repair bandwidth by contacting $d>k$ helpers and downloading only a fraction of data from each. Guruswami and Wootters first proposed a linear repair scheme for Reed-Solomon (RS) codes, showing that they can be repaired with lower bandwidth than the naive approach. The existence of RS codes achieving the MSR point (RS-MSR codes) nevertheless remained open until the breakthrough construction of Tamo, Barg, and Ye, which yields RS-MSR codes with subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$, where $p_{i}$ are distinct primes satisfying $p_{i}\equiv 1\pmod{s}$ and s=d+1ks=d+1-k.
In this paper, we present an improved construction of RS-MSR codes by eliminating the congruence condition $p_{i}\equiv 1\pmod{s}$. Consequently, our construction reduces the subpacketization by a multiplicative factor of $\phi(s)^{n}$ ( $\phi(\cdot)$ is Euler’s totient function) and broadens the range of feasible parameters for RS-MSR codes.
I Introduction
The exponential growth of data in distributed storage systems has necessitated the use of efficient erasure codes to ensure reliability. Maximum Distance Separable (MDS) codes, particularly Reed-Solomon (RS) codes, are preferred due to their optimal storage efficiency. However, a major drawback of traditional MDS codes is the high repair bandwidth required to recover a single failed node.
The concept of Regenerating Codes, introduced by Dimakis et al. [3], established the fundamental tradeoff between storage overhead and repair bandwidth. Minimum Storage Regenerating (MSR) codes achieve the cut-set bound for repair bandwidth while maintaining the MDS property.
Assume the data are encoded by a code $\mathcal{C}$ over a finite field $F$, where each code symbol is either a single element of $F$ or an $\ell$-dimensional vector over $F$. In the former case, $\mathcal{C}$ is called a scalar code; in the latter case, $\mathcal{C}$ is called an array code with subpacketization $\ell$. Most MSR constructions have focused on array codes, and many interesting results have been obtained, see [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. In contrast, constructing scalar MDS codes, particularly Reed–Solomon codes, that achieve the cut-set bound remained an open problem for several years. A landmark result by Tamo, Ye, and Barg [1] provided the first and, to date, only explicit construction of $(n,k)$ Reed–Solomon codes that are MSR codes, utilizing the linear repair scheme for scalar codes proposed by Guruswami and Wootters [2]. Their construction, referred to as RS-MSR codes, selects evaluation points from a field extension determined by a set of primes $p_{i}$ satisfying $p_{i}\equiv 1\pmod{s}$, where s=d+1ks=d+1-k and $d$ denotes the number of helper nodes participating in the repair of a failed node. The resulting subpacketization is $\ell=s\prod_{i=1}^{n}p_{i}$. A natural question is whether RS-MSR codes can achieve a smaller subpacketization. In this work, we present a follow-up to the construction in [1] to answer this question affirmatively.
Our main contribution is a basis-transformation technique composed of three parts: Euclidean Square Partition, Reshape, and Interference, which transforms a basis of the extension field into another basis that enables optimal repair of RS codes. Our results can be regarded as a generalization of [1] by relaxing the restriction on the choice of the field extension degrees $p_{i}$, namely, $p_{i}\equiv 1\pmod{s}$, thereby reducing the subpacketization by a factor of $\phi(s)^{n}$, where $\phi(\cdot)$ denotes Euler’s totient function.
The remainder of this paper is organized as follows. Section II provides the necessary preliminaries on RS codes and linear repair schemes. Section III presents our main results. Section IV concludes this paper.
II Preliminaries
Throughout this paper, we use the following notation unless stated otherwise.
  • For integers $a\leq b$, define $a,b]:=\{a,a+1,\dots,b\}$.
  • For a positive integer $n$, define $n]:=\{1,\dots,n\}$.
  • For any array $A$ of size $a\times b$, we index its entries by $(i,j)\in[0,a-1]\times[0,b-1]$, and denote by $A(i,j)$ the entry in the $i$-th row and $j$-th column.
II-A RS Codes and Repair Bandwidth
Let $E$ be a finite field. Let $\Omega=\{\omega_{1},\omega_{2},\ldots,\omega_{n}\}\subseteq E$ be a set of $n$ pairwise distinct elements, and let $v=(v_{1},\ldots,v_{n})\in(E^{*})^{n}$, where $E^{*}:=E\setminus\{0\}$. The generalized Reed–Solomon code $\operatorname{GRS}_{E}(n,k,\Omega,v)\subseteq E^{n}$ with evaluation set $\Omega$ and column multipliers $v$ is defined as the set
$\left\{\left(v_{1}f(\omega_{1}),\ldots,v_{n}f(\omega_{n})\right)\in E^{n}:\ f\in E[x],\ \deg f\leq k-1\right\}.$
If $v=(1,\ldots,1)$, then the GRS code is called a Reed–Solomon code and is denoted by $\operatorname{RS}_{E}(n,k,\Omega)$. It is well known (see [20]) that
$\left(\operatorname{RS}_{E}(n,k,\Omega)\right)^{\perp}=\operatorname{GRS}_{E}(n,n-k,\Omega,u),$
where $u_{i}=\prod_{j\neq i}(\omega_{i}-\omega_{j})^{-1}$ for $i=1,\ldots,n$.
Given an $(n,k)$ MDS array code $\mathcal{C}$ over a finite field $F$ with subpacketization $\ell$ (i.e., each node stores a vector $c_{i}\in F^{\ell}$). Suppose that the $i$-th node $c_{i}$ is erased, the repair bandwidth $N(\mathcal{C},i,\mathcal{R})$ of $c_{i}$ is defined as the total amount of information (measured in symbols over $F$) downloaded from a helper set $\mathcal{R}\subseteq[n]\setminus\{i\}$ with $|\mathcal{R}|=d$ to reconstruct $c_{i}$. The cut-set bound of [3] states that the repair bandwidth $N(\mathcal{C},i,\mathcal{R})$ satisfies
$N(\mathcal{C},i,\mathcal{R})\geq\frac{d\ell}{d-k+1}.$ (1)
We call $\mathcal{C}$ an MSR code if it achieves (1) with equality for all $i\in[n]$.
Assume that $E$ is an extension field of $F$ with $E:F]=\ell$, then we can regard an RS code over $E$ as an MDS array code over $F$ with subpacketization $\ell$. An RS code over $E$ is called an RS-MSR code if it is an MSR code when we regard it as an MDS array code over $F$.
II-B Linear Repair Schemes for scalar MDS codes
A key result in [2] is that the repair of a failed node $c_{i}$ can be performed by downloading the values of $\ell$ dual codewords at the helper nodes. Specifically, let $\mathcal{C}^{\perp}$ be the dual code of $\mathcal{C}$. If there exists a set of dual codewords $\{h_{1},\dots,h_{\ell}\}\subset\mathcal{C}^{\perp}$ such that their $i$-th coordinates span $E$ over $F$, then the node $c_{i}$ can be repaired.
Theorem 1 ([2]).
Let $\mathcal{C}\subset E^{n}$ be a scalar linear MDS code of length $n$. Let $F$ be a subfield of $E$ such that $E:F]=\ell$. For a given $i\in[n]$ the following statements are equivalent.
  1. (1)
    There is a linear repair scheme of the node $c_{i}$ over $F$ such that the repair bandwidth $N(\mathcal{C},i,\mathcal{R})\leq b$.
  2. (2)
    There is a subset of codewords $P_{i}\subseteq\mathcal{C}^{\perp}$ with size $|P_{i}|=\ell$ such that
    ${\rm dim}_{\mathbb{F}}(\{x_{i}:x\in P_{i}\})=\ell,$
    and
    $b\geq\sum_{j\in[n]\setminus\{i\}}{\rm dim}_{F}(\{x_{j}:x\in P_{i}\}).$
II-C Reed-Solomon codes as MSR codes
In this subsection, we introduce the main idea of the construction of RS-MSR codes in [1]. Precisely, given any positive integers $k<d\leq n-1$, the goal is to construct an $(n,k)$ RS code $\mathcal{C}$ achieving the cut-set bound for the repair of any single node from any $d$ helper nodes.
Let $\mathbb{F}_{p}$ be a finite field of prime order $p$. Denote s:=dk+1s:=d-k+1 and let $p_{1},\dots,p_{n}$ be $n$ distinct primes such that
$p_{i}\equiv 1\;\text{mod}\,s\;\;\text{~for all~}i=1,2,\dots,n.$ (2)
For $i=1,\dots,n$, let $\alpha_{i}$ be an element of an extension field of $\mathbb{F}_{p}$ such that $\mathbb{F}_{p}(\alpha_{i}):\mathbb{F}_{p}]=p_{i}$, and let
$\mathbb{F}:=\mathbb{F}_{p}(\alpha_{1},\dots,\alpha_{n}).$
Then it can be checked that $\mathbb{F}:\mathbb{F}_{p}]=\prod_{i=1}^{n}p_{i}$. Next, we define $n$ distinct subfields $F_{i}$ of the field $\mathbb{F}$ and one extension field $\mathbb{E}$ of $\mathbb{F}$.
  1. 1.
    For $i=1,\dots,n$, define $F_{i}=\mathbb{F}_{p}(\{\alpha_{j}:j\neq i\})$. Note that $\mathbb{F}=F_{i}(\alpha_{i})$ and $\mathbb{F}:F_{i}]=p_{i}$.
  2. 2.
    The field $\mathbb{E}$ is defined to be the degree-$s$ extension of the field $\mathbb{F}$, i.e., there exists an element $\beta\in\mathbb{E}$ of degree $s$ over $\mathbb{F}$ such that $\mathbb{E}=\mathbb{F}(\beta).$ We also have $\mathbb{E}:F_{i}]=sp_{i}$ for all $i$.
The construction of RS-MSR codes in [1] is as follows.
Theorem 2 ([1]).
Let $k,n,d$ be any positive integers such that $k<d<n.$ Let $\Omega=\{\alpha_{1},\dots,\alpha_{n}\}$, where $\alpha_{i},i=1,\dots,n$ is an element of degree $p_{i}$ over $\mathbb{F}_{p}$ and $p_{i}$ is the $i$-th smallest prime that satisfies (2). The code $\mathcal{C}:=\text{\rm RS}_{\mathbb{E}}(n,k,\Omega)$ achieves the cut-set bound for the repair of any single node from any $d$ helper nodes. In other words, $\mathcal{C}$ is an $(n,k)$ MSR code with repair degree $d$ and subpacketization $s\prod_{i=1}^{n}p_{i}$.
The proof of Theorem 2 relies on the following lemmas.
Lemma 1 ([21]).
Let $E$ be an extension field of finite field $F$ with $E:F]=\ell$. Assume $\{\zeta_{1},\dots,\zeta_{\ell}\}$ and $\{\mu_{1},\dots,\mu_{\ell}\}$ are dual basis of $E$ over $F$, i.e., ${\rm tr}(\zeta_{i}\mu_{j})=\delta_{ij}$, where ${\rm tr}(\cdot)$ denote the trace function from $E$ to $F$ and $\delta_{ij}$ is the Kronecker delta. Then for any $\gamma\in E$,
$\gamma=\sum_{i=1}^{\ell}{\rm tr}(\zeta_{i}\gamma)\mu_{i}.$
For a vector space $V$ over a field $F$ and a set of vectors $A=\left(a_{1},\ldots,a_{l}\right)\subset V$, let $\operatorname{Span}_{F}(A)=\left\{\sum_{i=1}^{l}\gamma_{i}a_{i},\gamma_{i}\in F\right\}$ be the span of $A$ over $F$.
Lemma 2 ([1]).
Let $\alpha_{i},\beta,F_{i}$ be as defined previously. Given $i\in[n]$, define the following vector spaces over $F_{i}$:
$S_{i}^{(1)}=\operatorname{Span}_{F_{i}}\big(\beta^{u}\alpha_{i}^{u+qs},u\in[0,s-1],q\in[0,\frac{p_{i}-1}{s}-1]\big)$
$S_{i}^{(2)}=\operatorname{Span}_{F_{i}}\Big(\sum_{t=0}^{s-1}\beta^{t}\alpha_{i}^{p_{i}-1}\Big)$
$S_{i}=S_{i}^{(1)}+S_{i}^{(2)}.$
Then
$\dim_{F_{i}}S_{i}=p_{i}\text{ and }S_{i}+S_{i}\alpha_{i}+\dots+S_{i}\alpha_{i}^{s-1}=\mathbb{E}.$ (3)
In what follows, we present a compressed version of the proof of Theorem 2 from [1] to highlight the role of Lemma 2 in determining the subpacketization.
Let $\mathcal{C}={\rm RS}_{\mathbb{E}}(n,k,\Omega)$, $\mathbb{F}$ be a subfield of $\mathbb{E}$ with $\mathbb{E}:\mathbb{F}]=\ell$. Let $c=(c_{1},c_{2},\dots,c_{n})$ be a codeword of $\mathcal{C}$, assume that $c_{i}$ is erased, it can be regenerated by the following procedure:
  • (1)
    By Lemma 2, there exists a vector space $S_{i}$ over the field $F_{i}$ such that
    $\dim_{F_{i}}S_{i}=p_{i}\text{ and }S_{i}+S_{i}\alpha_{i}+\dots+S_{i}\alpha_{i}^{s-1}=\mathbb{E},$
    let $e_{1},\dots,e_{p_{i}}$ be an arbitrary basis of the subspace $S_{i}$ over the field $F_{i}$.
  • (2)
    Choose $R_{i}\subset[n]\setminus\{i\}$ with $|R_{i}|=d$, where d=s+k1d=s+k-1.
  • (3)
    Let $h(x)=\prod_{j\in[n]\setminus(R_{i}\cup\{i\})}(x-\alpha_{j})$, then for $0\leq t\leq s-1$, $(v_{1}\alpha_{1}^{t}h(\alpha_{1}),\dots,v_{n}\alpha_{n}^{t}h(\alpha_{n}))\in C^{\perp},$ hence $\sum_{j=1}^{n}c_{j}v_{j}\alpha_{j}^{t}h(\alpha_{j})=0$
  • (4)
    Assume that $R_{i}=\{i_{1},\dots,i_{d}\}$, then
    $\sum_{j=1}^{d}c_{i_{j}}v_{i_{j}}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}})=0.$
  • (5)
    Let ${\rm tr}_{i}:={\rm tr}_{\mathbb{E}/F_{i}}$ denote the trace function from $\mathbb{E}$ to $F_{i}$. By Lemma 1, the set
    $\left\{{\rm tr}_{i}(c_{i}e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$
    can be used to recover $c_{i}$, as $\left\{e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$ forms a basis of $\mathbb{E}$ over $F_{i}$.
  • (6)
    For each $0\leq t\leq s-1$ and $1\leq m\leq p_{i}$,
    ${\rm tr}_{i}(c_{i}e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))$ $=-\sum_{j=1}^{d}{\rm tr}_{i}(c_{i_{j}}e_{m}v_{i_{j}}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}}))$
    $=-\sum_{j=1}^{d}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}}){\rm tr}_{i}(c_{i_{j}}e_{m}v_{i_{j}})$
    Hence, from each helper node, for example, the $i_{1}$-th node in $R_{i}$, one only need to calculate $\{{\rm tr}_{i}(c_{i_{1}}v_{i_{1}}e_{m})\}_{1\leq m\leq p_{i}}$, i.e. $p_{i}$ symbols form $F_{i}$. Consequently, the total repair bandwidth is $dp_{i}=\frac{sp_{i}d}{d+1-k}$, meet the cut-set bound (1).
Remark 1.
From the compressed proof of Theorem 2, the key to optimal repair is to find a $\mathbb{E}$-subspace $S_{i}$ satisfying (3). Lemma 2 provides such subspaces under the congruence condition $p_{i}\equiv 1\pmod{s}$, which yields an RS-MSR code with subpacketization $\ell=s(\prod_{i=1}^{n}p_{i}).$ In the next section, we show that this congruence condition can be removed, which further reduces the subpacketization.
III Main results
In this section, we generalize Lemma 2 by constructing an explicit $S$ satisfying (3) without assuming $p_{i}\equiv 1\pmod{s}$, thereby obtaining an RS–MSR code with smaller subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$ for any choice of distinct primes $p_{1},\ldots,p_{n}>s$.
Lemma 3.
Assume that $p$ and $s$ are positive integers such that $p>s$. Let $\beta$ be a generator of $\mathbb{K}$ over $\mathbb{E}$ with $\mathbb{K}:\mathbb{E}]=s$, and let $\alpha$ be a generator of $\mathbb{E}$ over $\mathbb{F}$ with $\mathbb{E}:\mathbb{F}]=p$. Then there exists an $\mathbb{F}$-subspace $S\subseteq\mathbb{K}$ with $\dim_{\mathbb{F}}(S)=p$ such that
$\mathbb{K}=\bigoplus_{u=0}^{s-1}\beta^{u}\mathbb{E}=\sum_{u=0}^{s-1}\alpha^{u}S.$
The main idea for finding the desired subspace $S$ is to view the problem as a change of $\mathbb{F}$-basis. Viewing $\mathbb{K}$ as an $\mathbb{F}$-vector space of dimension $sp$, one checks that
$\bigl\{\alpha^{i}(\alpha\beta)^{j}:i=0,\dots,p-1,\ j=0,\dots,s-1\bigr\}$
forms an $\mathbb{F}$-basis of $\mathbb{K}$. We arrange it as a $p\times s$ array $B$ (Figure 1). For $(i,j)\in[0,p-1]\times[0,s-1]$,
$B(i,j)=\alpha^{i}(\alpha\beta)^{j}.$ (4)
Thus, proving Lemma 3 reduces to the problem of finding another $\mathbb{F}$-basis that can be arranged as an $s\times p$ array (Figure 2) such that the $\mathbb{F}$-span of the first row equals $S$, and each subsequent row is obtained by multiplying the first row by a successive power of $\alpha$.
[…]
Figure 1: Original ($p\times s$)
$\rightarrow\,?$
[…]
Figure 2: Target ($s\times p$)
We achieve this basis change via a transformation procedure composed of: Euclidean Square Partition, Reshape, and Interference.
Euclidean Square Partition
Let $p>s$, and write the Euclidean algorithm as
$p$ $=a_{0}s+b_{1},$ (5)
$s$ $=a_{1}b_{1}+b_{2},$
$b_{1}$ $=a_{2}b_{2}+b_{3},$
$\ \vdots$
$b_{m-2}$ $=a_{m-1}b_{m-1}+b_{m},$
$b_{m-1}$ $=a_{m}b_{m},$
where $b_{m}=\gcd(p,s)$.
Given a $p\times s$ array $A$, we partition its entries into square blocks by partitioning its index set
$\mathcal{I}(p,s)\triangleq[0,p-1]\times[0,s-1]$
into disjoint squares via the following deterministic recursive procedure.
Definition 1 (Euclidean Square Partition).
Let $h,w$ be positive integers and let
$\mathcal{I}(h,w)\triangleq[0,h-1]\times[0,w-1].$
We define a partition of $\mathcal{I}(h,w)$ into disjoint squares via the following recursive procedure. Let $t:=\min\{h,w\}$. Tile $\mathcal{I}(h,w)$ by placing as many $t\times t$ squares as possible along the longer side (left-to-right if $w\geq h$, and top-to-bottom if $h>w$). This leaves a (possibly empty) residual rectangle, denoted by $\mathcal{I}(h^{\prime},w^{\prime})$; if it is nonempty, apply the same procedure recursively to $\mathcal{I}(h^{\prime},w^{\prime})$. The resulting collection of squares forms a disjoint partition of $\mathcal{I}(h,w)$, called the Euclidean square partition of $\mathcal{I}(h,w)$ and denoted by $\mathcal{P}(h,w)$. Each square in $\mathcal{P}(h,w)$ is called a Euclidean square.
This partition induces a partition of the array $A$ into square subarrays: for each $Q\in\mathcal{P}(h,w)$, let
$A[Q]\triangleq\bigl(A(i,j)\bigr)_{(i,j)\in Q}.$
With a slight abuse of notation, we use the same symbol $Q$ to refer to both the index square $Q\subseteq\mathcal{I}(h,w)$ and the associated square block $A[Q]$ whenever no confusion arises.
Applying the above procedure to $(h,w)=(p,s)$ yields $\mathcal{P}(p,s)$. Moreover, the side lengths of the squares in $\mathcal{P}(p,s)$ form the multiset
$\Big\{\underbrace{s,\dots,s}_{a_{0}},\underbrace{b_{1},\dots,b_{1}}_{a_{1}},\dots,\underbrace{b_{m},\dots,b_{m}}_{a_{m}}\Big\},$
as induced by the Euclidean chain (5).
Reshape
Given the Euclidean Square Partition $\mathcal{P}(p,s)$, we define a reshape map $\mathsf{Reshape}_{p,s}(\cdot)$ that converts a $p\times s$ array into an $s\times p$ array by (i) taking a overall transpose, and then (ii) performing a block-wise transpose (as specified in Definition 2).
Definition 2 (Reshape).
Let $B$ be a $p\times s$ array and let $B^{\top}$ be the $s\times p$ array defined by $B^{\top}(i,j)=B(j,i)$. For each Euclidean square $Q$ in $B$ with top-left corner index $(x,y)$ and side length $t$, let $Q^{\top}$ be the corresponding $t\times t$ square in $B^{\top}$ with top-left corner index $(y,x)$. Define $R=\mathsf{Reshape}_{p,s}(B)$ as the array obtained from $B^{\top}$ by transposing every square $Q^{\top}$, equivalently, for every such $Q$ and all $0\leq u,v\leq t-1$,
R(y+u,x+v)=B(x+u,y+v).R(y+u,\ x+v)=B(x+u,\ y+v). (6)
Interference
Firstly, we provide the definition of the interference operation.
Definition 3.
Let $m$ be a positive integer, and let $x,j_{1},j_{2},\dots,j_{m}$ be non-negative integers. Denote $J$ as the set of all distinct integers among $j_{1},j_{2},\dots,j_{m}$. For any symbols $\alpha,\beta$, consider the sequence $\{a_{i}\}_{i=1}^{m}$ defined by
$a_{i}=\alpha^{x+i}\beta^{j_{i}},\qquad i=1,\dots,m,$
i.e., the exponents of $\alpha$ are consecutive. We define the interference operator $\mathrm{Int}(\cdot)$ by mapping each term $a_{i}$ to
$\mathrm{Int}(a_{i}):=\overline{a}_{i}=\alpha^{x+i}\left(\sum_{j\in J}\beta^{j}\right)$
We call $\overline{a}_{i}$ the interference associated with $a_{i}$, and we call $\{\overline{a_{i}}\}_{i=1}^{m}$ the interference of the sequence $\{a_{i}\}_{i=1}^{m}$.
Example 1.
If $a_{1}=\alpha^{2}\beta^{2}$, $a_{2}=\alpha^{3}\beta^{4}$, and $a_{3}=\alpha^{4}\beta^{4}$, then the interference of the sequence $\{a_{1},a_{2},a_{3}\}$ is
$\left\{\alpha^{2}(\beta^{2}+\beta^{4}),\alpha^{3}(\beta^{2}+\beta^{4}),\alpha^{4}(\beta^{2}+\beta^{4})\right\}.$
Execute the interference operator $\mathrm{Int}(\cdot)$ column-wise on the reshaped array $R$, and denote the resulting $s\times p$ array by $\overline{R}:=\mathrm{Int}(R)$.
The interference operation in Definition 3 requires that, within each column, the exponents of $\alpha$ form a consecutive sequence. This property is guaranteed by the following lemma.
Lemma 4 (Consecutive $\alpha$-exponents after the reshape).
For every $(i,j)\in[0,s-1]\times[0,p-1]$, we have
$R(i,j)=\alpha^{i+j}\,\beta^{\tau(i,j)}$ (7)
for some integer $\tau(i,j)\in[0,s-1]$ depending on $(i,j)$. In particular, for each fixed column index $j$, the $\alpha$-exponents appearing in the $j$-th column of $R$ are exactly
j,j+1,,j+s1,j,\ j+1,\ $\dots$,\ j+s-1,
hence form a consecutive sequence of length $s$.
Proof.
By Definition 2, for any Euclidean square $Q\in\mathcal{P}(p,s)$ with top-left corner index $(x,y)$ and side length $t$,
$R(y+u,\ x+v)=B(x+u,\ y+v),\qquad 0\leq u,v\leq t-1.$
Substituting the expression of $B$ in (4) gives
$R(y+u,\ x+v)=\alpha^{x+u}(\alpha\beta)^{y+v}=\alpha^{x+u+y+v}\beta^{y+v}.$
Let $i=y+u$ and $j=x+v$. Then $x+u+y+v=i+j$, and hence
$R(i,j)=\alpha^{i+j}\beta^{y+v},$
which proves (7) with $\tau(i,j)=y+v\in[0,s-1]$. For fixed $j$, the exponent of $\alpha$ in $R(i,j)$ equals $i+j$, so as $i$ runs from 0 to s1s-1 we obtain the consecutive set $\{j,\dots,j+s-1\}$. ∎
Remark 2.
$\mathrm{Int}(\cdot)$ acts as the identity on the first $a_{0}s$ columns of $R$.
The entire transformation procedure is illustrated in Fig. 3 using a concrete example.
[…]
Figure 3: The transformation procedure for $(p,s)=(7,5)$.
Define
$S:=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(0,t):\,0\leq t\leq p-1\}.$ (8)
We claim that this subspace is exactly the desired $S$ in Lemma 3. Before proving this claim, we first give the explicit form of $S$.
Since $\mathrm{Int}(\cdot)$ acts as the identity on the first $a_{0}s$ columns of $R$, we decompose $S=S^{(1)}+S^{(2)},$ where
$S^{(1)}$ $=\operatorname{Span}_{\mathbb{E}}\{\overline{R}(0,qs+u):q\in[0,a_{0}-1],u\in[0,s-1]\},$
$S^{(2)}$ $=\operatorname{Span}_{\mathbb{E}}\{\overline{R}(0,a_{0}s+j):0\leq j\leq b_{1}-1\}.$
It is easy to verify that for $0\leq q\leq a_{0}-1$ and $0\leq u\leq s-1$,
$\overline{R}(0,qs+u)=R(0,qs+u)=\alpha^{qs}(\alpha\beta)^{u}.$
Hence $S^{(1)}$ is determined, and it suffices to determine $S^{(2)}$.
For $0\leq j\leq b_{1}-1$, define
$\mathcal{R}_{j}:=\overline{R}(0,a_{0}s+j),\\$
$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t}\quad(t\geq 1).$
By Definition 3, determining $\mathcal{R}_{j}$ amounts to identifying all $\beta$-exponents that appear in the $(a_{0}s+j)$-th column of $R$. We give an explicit expression for $\mathcal{R}_{j}$ and defer its proof to Appendix A.
Proposition 1.
For each $j\in\{0,1,\dots,b_{1}-1\}$,
$\mathcal{R}_{j}$ $=\alpha^{a_{0}s+j}\Bigg(\sum_{\begin{subarray}{c}1\leq t\leq\lceil m/2\rceil\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\end{subarray}}\ \sum_{q=0}^{a_{2t-1}-1}\beta^{\left(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\right)+q\,b_{2t-1}+j^{(2t-2)}}$
$\hskip 73.97733pt+\,\sum_{\begin{subarray}{c}1\leq t\leq\lfloor m/2\rfloor\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\\ \left\lfloor\frac{j^{(2t-2)}}{b_{2t}}\right\rfloor<a_{2t}\end{subarray}}\beta^{\left(\sum_{k=1}^{t}a_{2k-1}b_{2k-1}\right)+j^{(2t)}}\Bigg),$ (9)
Corollary 1.
When $m=1$, from Proposition 1 we have that for $j\in[0,b_{1}-1]$, $\mathcal{R}_{j}=\alpha^{a_{0}s}(\alpha\beta)^{j}\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}}.$
Proof.
We substitute $m=1$ into (1). Since $\lceil m/2\rceil=1$ and $\lfloor m/2\rfloor=0$, the second sum in (1) is empty, and the first sum contains only the term $t=1$. For $t=1$ we have $\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}=0$, hence the exponents contributed by the unique vertical term are
$tb_{1}+j,\qquad t=0,1,\dots,a_{1}-1.$
Therefore,
$\mathcal{R}_{j}=\alpha^{a_{0}s+j}\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}+j}=\alpha^{a_{0}s}(\alpha\beta)^{j}\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}}.$
Remark 3.
Lemma 2 corresponds to the special case of this corollary where $b_{1}=1$.
Corollary 2.
When $m=2$, from Proposition 1 we have that for $j\in[0,b_{1}-1]$,
$\mathcal{R}_{j}=\alpha^{a_{0}s}(\alpha\beta)^{j}\Bigg(\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}}+\beta^{a_{1}b_{1}-\left\lfloor j/b_{2}\right\rfloor b_{2}}\Bigg).$
Proof.
We substitute $m=2$ into (1). Here $\lceil m/2\rceil=1$ and $\lfloor m/2\rfloor=1$, hence both sums in (1) contain only the term $t=1$.
The vertical contribution (the first sum with $t=1$) is identical to the case $m=1$:
$\alpha^{a_{0}s+j}\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}+j}.$
The horizontal contribution: since $b_{1}=a_{2}b_{2}$ and $0\leq j\leq b_{1}-1$, we have $0\leq\lfloor j/b_{2}\rfloor\leq a_{2}-1$, so the condition $\big\lfloor j^{(0)}/b_{2}\big\rfloor<a_{2}$ holds for all such $j$. Therefore, the horizontal term contributes the exponent
$a_{1}b_{1}+j^{(2)}.$
Combining both contributions yields
$\mathcal{R}_{j}=\alpha^{a_{0}s+j}\Bigg(\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}+j}+\beta^{a_{1}b_{1}+j^{(2)}}\Bigg)=\alpha^{a_{0}s}(\alpha\beta)^{j}\Bigg(\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}}+\beta^{a_{1}b_{1}-\lfloor j/b_{2}\rfloor b_{2}}\Bigg),$
where we used $\beta^{a_{1}b_{1}+j^{(2)}}=\beta^{j}\beta^{a_{1}b_{1}-\lfloor j/b_{2}\rfloor b_{2}}$.
It is easy to verify that the example in Fig 3 coincides with this corollary.
Here we provide a sketch of proof of Lemma 3 and an illustrative example. The complete proof will be given in Appendix B.
The sketch of the proof of Lemma 3: from Remark 2, we only need to prove that the last $b_{1}$ columns of $R$ can be recovered by elements from vector space spanned from $\overline{R}$. Which can be done column by column and in each column the recovery order is from top to bottom. For example, in Fig. 3, the recovery proceeds as follows:
  1. 1.
    $R(0,5)=\overline{R}(0,5)-\overline{R}(3,2)-\overline{R}(1,4)$,
  2. 2.
    $R(1,5)=\overline{R}(1,5)-\overline{R}(4,2)-\overline{R}(2,4)$,
  3. 3.
    $R(2,5)=\overline{R}(2,5)-\alpha^{7}-\overline{R}(3,4)$,
  4. 4.
    $R(3,5)=\overline{R}(3,5)-\alpha^{8}-\overline{R}(4,4)$,
  5. 5.
    $R(4,5)=\overline{R}(4,5)-\alpha^{9}-\alpha^{9}\beta^{2}$,
  6. 6.
    $R(0,6)=\overline{R}(0,6)-\overline{R}(3,3)-\overline{R}(2,4)$,
  7. 7.
    $R(1,6)=\overline{R}(1,6)-\overline{R}(4,3)-\overline{R}(3,4)$,
  8. 8.
    $R(2,6)=\overline{R}(2,6)-\alpha^{8}\beta-\overline{R}(4,4)$,
  9. 9.
    $R(3,6)=\overline{R}(3,6)-\alpha^{9}\beta-R(4,5)$,
  10. 10.
    $R(4,6)=\overline{R}(4,6)-\alpha^{10}\beta-\alpha^{10}\beta^{3}$.
The above recovery is valid because
$\alpha^{7},\alpha^{8},\alpha^{9}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,0)\}_{t=0}^{4},R(0,5),R(1,5)\},$
$\alpha^{9}\beta^{2}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,2)\}_{t=0}^{4},R(2,5),R(3,5)\},$
$\alpha^{8}\beta,\alpha^{9}\beta,\alpha^{10}\beta$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,1)\}_{t=0}^{4},R(0,6),R(1,6)\},$
$\alpha^{10}\beta^{3}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,3)\}_{t=0}^{4},R(2,6),R(3,6)\},$
We are now ready to present our improved construction of RS–MSR codes. The optimal repair scheme is essentially the same as that in [1], see Section II-C.
Theorem 3.
Let $k,n,d$ be positive integers such that $k<d<n$, and set s:=d+1ks:=d+1-k. Given a finite field $\mathbb{F}_{q}$, let $p_{i}$ be the $i$-th smallest prime greater than $s$. Choose elements $\alpha_{1},\dots,\alpha_{n}$ such that $\deg_{\mathbb{F}_{q}}(\alpha_{i})=p_{i}$ for $i=1,\dots,n$, and set $\mathbb{F}:=\mathbb{F}_{q}(\alpha_{1},\dots,\alpha_{n})$. Let $\mathbb{E}$ be an extension field of $\mathbb{F}$ such that $\mathbb{E}:\mathbb{F}]=s$, and let $\Omega=\{\alpha_{1},\dots,\alpha_{n}\}\subseteq\mathbb{F}$. Then the code $\mathcal{C}:=\mathrm{RS}_{\mathbb{E}}(n,k,\Omega)$ is an $(n,k)$ MSR code over $\mathbb{F}_{q}$ with repair degree $d$ and subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$.
Remark 4 (The deduction of subpacketization).
Let $p_{i}$ denote the $i$-th prime greater than a constant $s$, and let $p_{i}^{\prime}$ denote the $i$-th prime satisfying $p_{i}^{\prime}\equiv 1\pmod{s}$. From the prime number theorem and the prime number theorem in arithmetic progressions [22], it can be shown that for $i\to\infty$,
$p_{i}\sim i\log i,~~p_{i}^{\prime}\sim\varphi(s)\,i\log i.$
Therefore,
$\frac{\ell_{\mathrm{TYB}}=\prod_{i=1}^{n}p_{i}^{\prime}}{\ell_{\mathrm{new}}=\prod_{i=1}^{n}p_{i}}\sim\varphi(s)^{n}.$
IV Conclusion
In this paper, we improve the construction of RS-MSR codes by showing that the previously imposed congruence condition on the primes $p_{i}$ is unnecessary, thereby reducing the subpacketization. A natural direction for future work is to determine whether the subpacketization can be reduced further, or to establish new lower bounds proving the tightness of our result.
Appendix A Proof of Proposition 1
Recall that determining $\mathcal{R}_{j}$ amounts to identifying the $\beta$-exponents appearing in the $(a_{0}s+j)$-th column of $R$, where $j\in\{0,1,\dots,b_{1}-1\}$. For simplicity, let $R_{0}$ denote the last $b_{1}$ columns of $R$. The following useful fact follows directly from the definition of the map $\mathsf{Reshape}_{p,s}(\cdot)$.
Fact: After the $\mathsf{Reshape}_{p,s}(\cdot)$, the squares in $R$ corresponding to Euclidean squares in $B$ form an Euclidean square partition of $R$, so does the squares in $R_{0}$, hence we can denote the set of all squares in $R_{0}$ by $\mathcal{P}(s,b_{1})$.
From the proof of Lemma 4, the exponents of $\beta$ inside each column of any Euclidean square $P$ in $R_{0}$ are the same. This motivates the following definition..
Definition 4 (Intersection indices $(y,v)$ for a fixed column).
Fix $j\in\{0,1,\dots,b_{1}-1\}$. Define $\mathcal{YV}(j)$ to be the set of all pairs $(y,v)$ for which there exist a Euclidean square $Q\in\mathcal{P}(s,b_{1})$ with upper-left corner $(y,x)$ and side length $t$ such that
$j\in\{x,x+1,\dots,x+t-1\}\quad\text{and}\quad v:=j-x\in\{0,1,\dots,t-1\}.$
Equivalently, $(y,v)\in\mathcal{YV}(j)$ if and only if the column indexed by $j$ in $\mathcal{I}(s,b_{1})$ intersects a Euclidean square whose upper-left row coordinate is $y$ with within-square column index $v$.
From Lemma 4, it follows immediately that
Proposition 2 (Reduction of $\mathcal{R}_{j}$ to $\mathcal{YV}(j)$).
For each $j\in\{0,1,\dots,b_{1}-1\}$, we have
$\mathcal{R}_{j}=\alpha^{a_{0}s+j}\sum_{(y,v)\in\mathcal{YV}(j)}\beta^{\,y+v}.$ (10)
Hence the core task is to characterize $\mathcal{YV}(j)$ for each $j$. This is purely a question about the Euclidean square partition $\mathcal{P}(s,b_{1})$. Set $b_{0}:=s$, the recursive construction of $\mathcal{P}(s,b_{1})$ begins with $b_{0}=a_{1}b_{1}+b_{2}$ and alternates between two types of decomposition steps:
  • the (2t1)(2t\!-\!1)-st vertical decomposition step, induced by
    $b_{2t-2}=a_{2t-1}b_{2t-1}+b_{2t},$ (11)
    which places $a_{2t-1}$ squares of side length $b_{2t-1}$ along the vertical direction;
  • the $(2t)$-th horizontal decomposition step, induced by
    $b_{2t-1}=a_{2t}b_{2t}+b_{2t+1},$ (12)
    which places $a_{2t}$ squares of side length $b_{2t}$ along the horizontal direction.
These steps are applied to the successive residual rectangles produced in the recursion.
Fix $j\in\{0,1,\dots,b_{1}-1\}$ and track the fixed global column $a_{0}s+j$ through the Euclidean recursion of $\mathcal{P}(s,b_{1})$. At any stage of the recursion, we describe the position of this fixed global column by its local column coordinate within the current residual rectangle. Vertical decomposition steps only cut the residual rectangle along the vertical direction and therefore preserve its horizontal interval, hence they do not change the local column coordinate. In contrast, at the $(2t)$-th horizontal decomposition step induced by $b_{2t-1}=a_{2t}b_{2t}+b_{2t+1}$, the current residual rectangle of width $b_{2t-1}$ is partitioned, from left to right, into $a_{2t}$ squares of side length $b_{2t}$ and one residual sub-rectangle of width $b_{2t+1}$. Accordingly, we define the local column coordinate recursively by
$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t}\quad(t\geq 1),$ (13)
so that $j^{(2t)}$ is the remainder of $j^{(2t-2)}$ upon division by $b_{2t}$.
Key observation: the column $a_{0}s+j$ enters the (2t1)(2t\!-\!1)-st vertical decomposition step if and only if, at each preceding horizontal decomposition step $(2k)$ with $k\in\{1,\dots,t-1\}$, the column lies in the residual sub-rectangle (rather than in one of the $a_{2k}$ squares of side length $b_{2k}$), i.e.,
$\forall k\in\{1,\dots,t-1\}:\ \Big\lfloor\frac{j^{(2k-2)}}{b_{2k}}\Big\rfloor=a_{2k}.$ (14)
Then we have:
  • (Vertical case: squares produced at the (2t1)(2t\!-\!1)-st vertical step). At the (2t1)(2t\!-\!1)-st vertical step, the recursion places $a_{2t-1}$ squares of side length $b_{2t-1}$ stacked vertically. Since these squares cover the full current width $b_{2t-1}$, the column intersects all of them, with the same within-square column index $v=j^{(2t-2)}$. Their upper-left row coordinates are
    $y=\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\qquad q=0,1,\dots,a_{2t-1}-1.$
    Hence $\mathcal{YV}(j)$ contains the $a_{2t-1}$ pairs
    $(y,v)=\Bigg(\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\ j^{(2t-2)}\Bigg),\qquad q=0,1,\dots,a_{2t-1}-1.$
  • (Horizontal terminating case: the unique square produced at the $(2t)$-th horizontal step). Now consider the subsequent $(2t)$-th horizontal step. If
    $\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor<a_{2t},$
    then the column lies in one of the $a_{2t}$ squares of side length $b_{2t}$ produced in this step, hence intersects exactly one such square. Its upper-left row coordinate equals
    $y=\sum_{k=1}^{t}a_{2k-1}b_{2k-1},$
    and the within-square column index is $v=j^{(2t)}$. Thus $\mathcal{YV}(j)$ contains the additional (terminating) pair
    $(y,v)=\Bigg(\sum_{k=1}^{t}a_{2k-1}b_{2k-1},\ j^{(2t)}\Bigg).$
    If instead $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor=a_{2t}$, then the column lies in the residual sub-rectangle and this horizontal step contributes no pair, as the recursion continues to $(2t+1)$-th vertical decomposition.
Collecting the contributions from the vertical case (for all reachable depths $t$) and from the horizontal terminating case (whenever termination occurs at step $(2t)$) yields the following closed form:
$\mathcal{R}_{j}$ $=\alpha^{a_{0}s+j}\Bigg(\sum_{\begin{subarray}{c}1\leq t\leq\lceil m/2\rceil\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\end{subarray}}\ \sum_{q=0}^{a_{2t-1}-1}\beta^{\left(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\right)+q\,b_{2t-1}+j^{(2t-2)}}$
$\hskip 73.97733pt+\,\sum_{\begin{subarray}{c}1\leq t\leq\lfloor m/2\rfloor\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\\ \left\lfloor\frac{j^{(2t-2)}}{b_{2t}}\right\rfloor<a_{2t}\end{subarray}}\beta^{\left(\sum_{k=1}^{t}a_{2k-1}b_{2k-1}\right)+j^{(2t)}}\Bigg),$
which finished the proof of Proposition 1.
Appendix B Proof of Lemma 3
Let
$K=\sum_{u=0}^{s-1}\alpha^{u}S=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(u,v):(u,v)\in\mathcal{I}(s,p)\}.$
From previous analysis, we only need to prove that for $(i,j)\in\mathcal{I}(s,b_{1})$, $R(i,a_{0}s+j)\in K$.
Fix $(i,j)\in\mathcal{I}(s,b_{1})$, let $\tau(\cdot,\cdot)$ be as in Lemma 4, then
$R(i,j):=\alpha^{a_{0}s+i+j}\beta^{\tau(i,a_{0}s+j)}.$
Equivalently, if $(i,j)$ lies in the unique Euclidean square whose upper-left corner is $(y,x)$, then Lemma 4 gives $\tau(i,a_{0}s+j)=y+(j-x)=j+(y-x)$, hence
$R(i,j)=\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{\,y-x}.$
Since $\mathcal{R}_{j}\in S$, for every $i\in[0,s-1]$ we have $\alpha^{i}\mathcal{R}_{j}\in K$. By Proposition 1 together with Lemma  4, the element $\alpha^{i}\mathcal{R}_{j}$ has the form
$\alpha^{i}\mathcal{R}_{j}=\alpha^{a_{0}s+i}(\alpha\beta)^{j}\Big(\beta^{y-x}+\sum_{\begin{subarray}{c}Q\text{ intersects col }j\\ Q\neq Q_{0}\end{subarray}}\beta^{y_{Q}-x_{Q}}\Big),$
where $Q_{0}$ is the Euclidean square containing $(i,j)$ and, for each other intersected square $Q$, $(y_{Q},x_{Q})$ denotes its upper-left corner. Consequently,
$R(i,j)=\alpha^{i}\mathcal{R}_{j}-\alpha^{a_{0}s+i}(\alpha\beta)^{j}\sum_{\begin{subarray}{c}Q\text{ intersects col }j\\ Q\neq Q_{0}\end{subarray}}\beta^{y_{Q}-x_{Q}}.$ (15)
Therefore it is enough to show that for every $Q\neq Q_{0}$ intersecting column $j$,
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}\in K.$ (16)
We prove (16) by induction on the column index $j$ ordered by the sequence of quotients appearing in the Euclidean recursion. Let $j^{(0)}:=j$ and $j^{(2t)}$ be defined by (13). Order columns by the lexicographic order of the finite sequence
$\Big(\Big\lfloor\frac{j^{(0)}}{b_{2}}\Big\rfloor,\ \Big\lfloor\frac{j^{(2)}}{b_{4}}\Big\rfloor,\ \Big\lfloor\frac{j^{(4)}}{b_{6}}\Big\rfloor,\ \ldots\Big),$
where the sequence stops when the reachability condition 14 first fails. Assume that for every column $j^{\prime}$ preceding $j$ in this order and every $i^{\prime}\in[0,s-1]$, we already have $R(i^{\prime},j^{\prime})\in K$.
Fix an intersected square $Q\neq Q_{0}$ in column $j$. If $Q$ is a terminating square arising from a horizontal decomposition step, then its upper-left column coordinate is of the form $x_{Q}\equiv 0\pmod{b_{2t}}$ for the corresponding step size $b_{2t}$, and the contribution of $Q$ in $\alpha^{i}\mathcal{R}_{j}$ is the term $\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}$ (this is exactly the terminating term in (1)). Choose an integer $z\geq 1$ such that $(i+zb_{2t},\,j-zb_{2t})\in\mathcal{I}(s,b_{1})$ lies in the same terminating square pattern at that step (such a $z$ exists because the terminating square has side length $b_{2t}$ and column $j$ lies in a block with $\lfloor j^{(2t-2)}/b_{2t}\rfloor>0$ at that step). Then, by the definition of $r(\cdot,\cdot)$ via Lemma 4 and the identity
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=\alpha^{a_{0}s+i+zb_{2t}}(\alpha\beta)^{j-zb_{2t}}\beta^{y_{Q}-(x_{Q}-zb_{2t})}=\alpha^{a_{0}s+(i+zb_{2t})+(j-zb_{2t})}\beta^{\tau(i+zb_{2t},a_{0}s+j-zb_{2t})},$
we obtain
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=R(i+zb_{2t},\,j-zb_{2t}).$
Moreover, $j-zb_{2t}$ precedes $j$ in the above lexicographic order because the quotient $\lfloor j^{(2t-2)}/b_{2t}\rfloor$ strictly decreases while all earlier quotients remain unchanged. Hence, by the induction hypothesis, $R(i+zb_{2t},\,j-zb_{2t})\in K$, which proves (16) for terminating squares.
If $Q$ is a vertical square in the Euclidean partition, then its contribution in (1) belongs to one of the vertical progressions at some vertical step, and the corresponding exponent difference $y_{Q}-x_{Q}$ differs from the target difference yxy-x by a positive multiple of the vertical step size at that stage. In this case, the element $\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}$ can be rewritten as
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=(\alpha\beta)^{j+(y_{Q}-x_{Q})}\alpha^{a_{0}s+i-(y_{Q}-x_{Q})}.$
Since $Q\neq Q_{0}$ and $(i,j)\in Q_{0}$, the row index $i$ lies strictly above the vertical block indexed by $Q$ at that step, hence $0\leq a_{0}s+i-(y_{Q}-x_{Q})\leq a_{0}s-1$. Therefore $(\alpha\beta)^{j+(y_{Q}-x_{Q})}\alpha^{a_{0}s+i-(y_{Q}-x_{Q})}\in K$ by $\sum_{i=0}^{s-1}\alpha^{i}S^{(1)}\in K$, proving (16) for vertical squares.
Thus every term in the sum on the right-hand side of (15) belongs to $K$, so (15) implies $R(i,j)\in K$. By induction over the above column order, we have $R(i,j)\in K$ for all $(i,j)\in\mathcal{I}(s,b_{1})$, which finished the proof.

### v2

@@
 Lemma 1 ([21]).
-Let $E$ be an extension field of finite field $F$ with $E:F]=\ell$.
-Assume $\{\zeta_{1},\dots,\zeta_{\ell}\}$ and $\{\mu_{1},\dots,\mu_{\ell}\}$ are dual basis of $E$ over $F$, i.e., ${\rm tr}(\zeta_{i}\mu_{j})=\delta_{ij}$, where ${\rm tr}(\cdot)$ denote the trace function from $E$ to $F$ and $\delta_{ij}$ is the Kronecker delta.
+Let $E$ be an extension field of a finite field $F$ with $E:F]=\ell$.
+Assume $\{\zeta_{1},\dots,\zeta_{\ell}\}$ and $\{\mu_{1},\dots,\mu_{\ell}\}$ are dual bases of $E$ over $F$, i.e., ${\rm tr}(\zeta_{i}\mu_{j})=\delta_{ij}$, where ${\rm tr}(\cdot)$ denotes the trace function from $E$ to $F$ and $\delta_{ij}$ is the Kronecker delta.
 Then for any $\gamma\in E$,
 $\gamma=\sum_{i=1}^{\ell}{\rm tr}(\zeta_{i}\gamma)\mu_{i}.$
-For a vector space $V$ over a field $F$ and a set of vectors $A=\left(a_{1},\ldots,a_{l}\right)\subset V$, let $\operatorname{Span}_{F}(A)=\left\{\sum_{i=1}^{l}\gamma_{i}a_{i},\gamma_{i}\in F\right\}$ be the span of $A$ over $F$.
+For a vector space $V$ over a field $F$ and a set of vectors $A=\left(a_{1},\ldots,a_{\ell}\right)\subset V$, define
+$\operatorname{Span}_{F}(A)=\left\{\sum_{i=1}^{\ell}\gamma_{i}a_{i}:\ \gamma_{i}\in F\right\}.$
 Lemma 2 ([1]).
@@
   • (3)
-    Let $h(x)=\prod_{j\in[n]\setminus(R_{i}\cup\{i\})}(x-\alpha_{j})$, then for $0\leq t\leq s-1$, $(v_{1}\alpha_{1}^{t}h(\alpha_{1}),\dots,v_{n}\alpha_{n}^{t}h(\alpha_{n}))\in C^{\perp},$ hence $\sum_{j=1}^{n}c_{j}v_{j}\alpha_{j}^{t}h(\alpha_{j})=0$
+    Let $h(x)=\prod_{j\in[n]\setminus(R_{i}\cup\{i\})}(x-\alpha_{j})$, then for $0\leq t\leq s-1$, $(v_{1}\alpha_{1}^{t}h(\alpha_{1}),\dots,v_{n}\alpha_{n}^{t}h(\alpha_{n}))\in\mathcal{C}^{\perp},$ hence $\sum_{j=1}^{n}c_{j}v_{j}\alpha_{j}^{t}h(\alpha_{j})=0$
   • (4)
-    Assume that $R_{i}=\{i_{1},\dots,i_{d}\}$, then
-    $\sum_{j=1}^{d}c_{i_{j}}v_{i_{j}}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}})=0.$
+    Assume that $R_{i}=\{i_{1},\dots,i_{d}\}$. Then for each $0\leq t\leq s-1$,
+    $c_{i}v_{i}\alpha_{i}^{t}h(\alpha_{i})+\sum_{j=1}^{d}c_{i_{j}}v_{i_{j}}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}})=0.$
   • (5)
@@
     $\left\{{\rm tr}_{i}(c_{i}e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$
-    can be used to recover $c_{i}$, as $\left\{e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$ forms a basis of $\mathbb{E}$ over $F_{i}$.
+    can be used to recover $c_{i}$, as $\left\{e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i})\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$ forms a basis of $\mathbb{E}$ over $F_{i}$.
   • (6)
@@
     $=-\sum_{j=1}^{d}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}}){\rm tr}_{i}(c_{i_{j}}e_{m}v_{i_{j}})$
-    Hence, from each helper node, for example, the $i_{1}$-th node in $R_{i}$, one only need to calculate $\{{\rm tr}_{i}(c_{i_{1}}v_{i_{1}}e_{m})\}_{1\leq m\leq p_{i}}$, i.e.
-$p_{i}$ symbols form $F_{i}$.
-Consequently, the total repair bandwidth is $dp_{i}=\frac{sp_{i}d}{d+1-k}$, meet the cut-set bound (1).
+    Hence, from each helper node (for example, the $i_{1}$-th node in $R_{i}$), one only needs to compute $\{{\rm tr}_{i}(c_{i_{1}}v_{i_{1}}e_{m})\}_{1\leq m\leq p_{i}}$, i.e., $p_{i}$ symbols from $F_{i}$.
+Consequently, the total repair bandwidth is
+    $dp_{i}=\frac{sp_{i}d}{d+1-k},$
+    which meets the cut-set bound (1).
 Remark 1.
@@
 III Main results
-In this section, we generalize Lemma 2 by constructing an explicit $S$ satisfying (3) without assuming $p_{i}\equiv 1\pmod{s}$, thereby obtaining an RS–MSR code with smaller subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$ for any choice of distinct primes $p_{1},\ldots,p_{n}>s$.
-Lemma 3.
+In this section, we present our main results, which show that the congruence condition $p_{i}\equiv 1\pmod{s}$ in Theorem 2 can be removed.
+As a result, we obtain an RS–MSR code with smaller subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$ for any choice of distinct primes $p_{1},\ldots,p_{n}>s$.
+From the discussion in Section II, the key is to show the existence of an $F_{i}$-subspace $S_{i}\subseteq\mathbb{E}$ that satisfies (3) without requiring the condition $p_{i}\equiv 1\pmod{s}$.
+Equivalently,
+Lemma 3 (Existence of an $\mathbb{F}$-subspace with $\alpha$-layer decomposition).
 Assume that $p$ and $s$ are positive integers such that $p>s$.
@@
 […]
-Figure 1: Original ($p\times s$)
+Figure 1: Original $p\times s$ array
 $\rightarrow\,?$
 […]
-Figure 2: Target ($s\times p$)
+Figure 2: Target $s\times p$ array
 We achieve this basis change via a transformation procedure composed of: Euclidean Square Partition, Reshape, and Interference.
@@
 where $b_{m}=\gcd(p,s)$.
-Given a $p\times s$ array $A$, we partition its entries into square blocks by partitioning its index set
-$\mathcal{I}(p,s)\triangleq[0,p-1]\times[0,s-1]$
-into disjoint squares via the following deterministic recursive procedure.
+We define a deterministic partition of any array of size $p\times s$ by partitioning its index set $\mathcal{I}(p,s)\triangleq[0,p-1]\times[0,s-1]$ into disjoint square blocks via the following recursive procedure.
 Definition 1 (Euclidean Square Partition).
-Let $h,w$ be positive integers and let
-$\mathcal{I}(h,w)\triangleq[0,h-1]\times[0,w-1].$
-We define a partition of $\mathcal{I}(h,w)$ into disjoint squares via the following recursive procedure.
+Let $h,w$ be positive integers and consider the rectangle $\mathcal{I}(h,w)=[0,h-1]\times[0,w-1]$.
 Let $t:=\min\{h,w\}$.
 Tile $\mathcal{I}(h,w)$ by placing as many $t\times t$ squares as possible along the longer side (left-to-right if $w\geq h$, and top-to-bottom if $h>w$).
-This leaves a (possibly empty) residual rectangle, denoted by $\mathcal{I}(h^{\prime},w^{\prime})$; if it is nonempty, apply the same procedure recursively to $\mathcal{I}(h^{\prime},w^{\prime})$.
-The resulting collection of squares forms a disjoint partition of $\mathcal{I}(h,w)$, called the Euclidean square partition of $\mathcal{I}(h,w)$ and denoted by $\mathcal{P}(h,w)$.
+This leaves a (possibly empty) residual rectangle $\mathcal{I}(h^{\prime},w^{\prime})$; if it is nonempty, apply the same procedure recursively to $\mathcal{I}(h^{\prime},w^{\prime})$.
+The resulting collection of squares forms a disjoint partition of $\mathcal{I}(h,w)$, called the Euclidean square partition of $\mathcal{I}(h,w)$, and is denoted as $\mathcal{P}(h,w)$.
 Each square in $\mathcal{P}(h,w)$ is called a Euclidean square.
-This partition induces a partition of the array $A$ into square subarrays: for each $Q\in\mathcal{P}(h,w)$, let
-$A[Q]\triangleq\bigl(A(i,j)\bigr)_{(i,j)\in Q}.$
-With a slight abuse of notation, we use the same symbol $Q$ to refer to both the index square $Q\subseteq\mathcal{I}(h,w)$ and the associated square block $A[Q]$ whenever no confusion arises.
 Applying the above procedure to $(h,w)=(p,s)$ yields $\mathcal{P}(p,s)$. Moreover, the side lengths of the squares in $\mathcal{P}(p,s)$ form the multiset
@@
 The interference operation in Definition 3 requires that, within each column, the exponents of $\alpha$ form a consecutive sequence. This property is guaranteed by the following lemma.
-Lemma 4 (Consecutive $\alpha$-exponents after the reshape).
+Lemma 4 (Consecutive $\alpha$-exponents in each reshaped column).
 For every $(i,j)\in[0,s-1]\times[0,p-1]$, we have
@@
 Since $\mathrm{Int}(\cdot)$ acts as the identity on the first $a_{0}s$ columns of $R$, we decompose $S=S^{(1)}+S^{(2)},$ where
-$S^{(1)}$ $=\operatorname{Span}_{\mathbb{E}}\{\overline{R}(0,qs+u):q\in[0,a_{0}-1],u\in[0,s-1]\},$
-$S^{(2)}$ $=\operatorname{Span}_{\mathbb{E}}\{\overline{R}(0,a_{0}s+j):0\leq j\leq b_{1}-1\}.$
+$S^{(1)}$ $=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(0,qs+u):q\in[0,a_{0}-1],u\in[0,s-1]\},$
+$S^{(2)}$ $=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(0,a_{0}s+j):0\leq j\leq b_{1}-1\}.$
 It is easy to verify that for $0\leq q\leq a_{0}-1$ and $0\leq u\leq s-1$,
@@
 For $0\leq j\leq b_{1}-1$, define
-$\mathcal{R}_{j}:=\overline{R}(0,a_{0}s+j),\\$
+$\mathcal{R}_{j}:=\overline{R}(0,a_{0}s+j),$
 $j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t}\quad(t\geq 1).$
@@
 The above recovery is valid because
-$\alpha^{7},\alpha^{8},\alpha^{9}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,0)\}_{t=0}^{4},R(0,5),R(1,5)\},$
-$\alpha^{9}\beta^{2}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,2)\}_{t=0}^{4},R(2,5),R(3,5)\},$
-$\alpha^{8}\beta,\alpha^{9}\beta,\alpha^{10}\beta$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,1)\}_{t=0}^{4},R(0,6),R(1,6)\},$
-$\alpha^{10}\beta^{3}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,3)\}_{t=0}^{4},R(2,6),R(3,6)\},$
+$\alpha^{7},\alpha^{8},\alpha^{9}$ $\in\operatorname{Span}_{\mathbb{F}}\{\{\overline{R}(t,0)\}_{t=0}^{4},R(0,5),R(1,5)\},$
+$\alpha^{9}\beta^{2}$ $\in\operatorname{Span}_{\mathbb{F}}\{\{\overline{R}(t,2)\}_{t=0}^{4},R(2,5),R(3,5)\},$
+$\alpha^{8}\beta,\alpha^{9}\beta,\alpha^{10}\beta$ $\in\operatorname{Span}_{\mathbb{F}}\{\{\overline{R}(t,1)\}_{t=0}^{4},R(0,6),R(1,6)\},$
+$\alpha^{10}\beta^{3}$ $\in\operatorname{Span}_{\mathbb{F}}\{\{\overline{R}(t,3)\}_{t=0}^{4},R(2,6),R(3,6)\}.$
 We are now ready to present our improved construction of RS–MSR codes. The optimal repair scheme is essentially the same as that in [1], see Section II-C.
@@
 Appendix A Proof of Proposition 1
-Recall that determining $\mathcal{R}_{j}$ amounts to identifying the $\beta$-exponents appearing in the $(a_{0}s+j)$-th column of $R$, where $j\in\{0,1,\dots,b_{1}-1\}$.
-For simplicity, let $R_{0}$ denote the last $b_{1}$ columns of $R$.
-The following useful fact follows directly from the definition of the map $\mathsf{Reshape}_{p,s}(\cdot)$.
-Fact: After the $\mathsf{Reshape}_{p,s}(\cdot)$, the squares in $R$ corresponding to Euclidean squares in $B$ form an Euclidean square partition of $R$, so does the squares in $R_{0}$, hence we can denote the set of all squares in $R_{0}$ by $\mathcal{P}(s,b_{1})$.
-From the proof of Lemma 4, the exponents of $\beta$ inside each column of any Euclidean square $P$ in $R_{0}$ are the same. This motivates the following definition..
+Recall that determining $\mathcal{R}_{j}$ amounts to identifying all $\beta$-exponents appearing in column $a_{0}s+j$ of the reshaped array $R$, where $j\in\{0,1,\dots,b_{1}-1\}$.
+Denote by $R_{\mathrm{1}}$ the subarray of $R$ consisting of its last $b_{1}$ columns.
+By Definition 2, the map $\mathsf{Reshape}_{p,s}(\cdot)$ preserves the Euclidean square partition structure; consequently, the squares in $R_{\mathrm{1}}$ induced by $\mathcal{P}(p,s)$ form precisely the Euclidean square partition $\mathcal{P}(s,b_{1})$ of the index set $\mathcal{I}(s,b_{1}):=[0,s-1]\times[0,b_{1}-1]$.
+From Lemma 4, the $\beta$-exponent is constant along each column of any Euclidean square in $R_{\mathrm{1}}$. This structural property leads to the following characterization.
 Definition 4 (Intersection indices $(y,v)$ for a fixed column).
 Fix $j\in\{0,1,\dots,b_{1}-1\}$.
-Define $\mathcal{YV}(j)$ to be the set of all pairs $(y,v)$ for which there exist a Euclidean square $Q\in\mathcal{P}(s,b_{1})$ with upper-left corner $(y,x)$ and side length $t$ such that
+Define $\mathcal{YV}(j)$ as the set of pairs $(y,v)$ for which there exists a Euclidean square $Q\in\mathcal{P}(s,b_{1})$ with top-left corner $(y,x)$ and side length $t$ satisfying
 $j\in\{x,x+1,\dots,x+t-1\}\quad\text{and}\quad v:=j-x\in\{0,1,\dots,t-1\}.$
-Equivalently, $(y,v)\in\mathcal{YV}(j)$ if and only if the column indexed by $j$ in $\mathcal{I}(s,b_{1})$ intersects a Euclidean square whose upper-left row coordinate is $y$ with within-square column index $v$.
-From Lemma 4, it follows immediately that
+Equivalently, $(y,v)\in\mathcal{YV}(j)$ if and only if column $j$ of $\mathcal{I}(s,b_{1})$ intersects a square whose top-left row coordinate is $y$ at local column index $v$ within that square.
+By Lemma 4, we obtain the following reduction.
 Proposition 2 (Reduction of $\mathcal{R}_{j}$ to $\mathcal{YV}(j)$).
-For each $j\in\{0,1,\dots,b_{1}-1\}$, we have
+For each $j\in\{0,1,\dots,b_{1}-1\}$,
 $\mathcal{R}_{j}=\alpha^{a_{0}s+j}\sum_{(y,v)\in\mathcal{YV}(j)}\beta^{\,y+v}.$ (10)
-Hence the core task is to characterize $\mathcal{YV}(j)$ for each $j$.
-This is purely a question about the Euclidean square partition $\mathcal{P}(s,b_{1})$.
-Set $b_{0}:=s$, the recursive construction of $\mathcal{P}(s,b_{1})$ begins with $b_{0}=a_{1}b_{1}+b_{2}$ and alternates between two types of decomposition steps:
+Thus the problem reduces to a combinatorial characterization of $\mathcal{YV}(j)$. Let $b_{0}:=s$ and consider the Euclidean algorithm applied to $(b_{0},b_{1})$:
+$b_{0}=a_{1}b_{1}+b_{2},\quad b_{1}=a_{2}b_{2}+b_{3},\quad\dots,\quad b_{m-2}=a_{m-1}b_{m-1}+b_{m},\quad b_{m-1}=a_{m}b_{m},$ (11)
+where $b_{m}=\gcd(s,b_{1})$ and $m\geq 1$ denotes the total number of division steps. The partition $\mathcal{P}(s,b_{1})$ is constructed recursively by alternating between two types of decomposition:
   • The (2t1)(2t\!-\!1)-st vertical decomposition step, induced by
     $b_{2t-2}=a_{2t-1}b_{2t-1}+b_{2t},$ (12)
-    which places $a_{2t-1}$ squares of side length $b_{2t-1}$ along the vertical direction;
+    which tiles the current residual rectangle of height $b_{2t-2}$ with $a_{2t-1}$ squares of side length $b_{2t-1}$ stacked vertically, leaving a residual rectangle of height $b_{2t}$.
   • The $(2t)$-th horizontal decomposition step, induced by
     $b_{2t-1}=a_{2t}b_{2t}+b_{2t+1},$ (13)
-    which places $a_{2t}$ squares of side length $b_{2t}$ along the horizontal direction.
-These steps are applied to the successive residual rectangles produced in the recursion.
-Fix $j\in\{0,1,\dots,b_{1}-1\}$ and track the fixed global column $a_{0}s+j$ through the Euclidean recursion of $\mathcal{P}(s,b_{1})$.
-At any stage of the recursion, we describe the position of this fixed global column by its local column coordinate within the current residual rectangle.
-Vertical decomposition steps only cut the residual rectangle along the vertical direction and therefore preserve its horizontal interval, hence they do not change the local column coordinate.
-In contrast, at the $(2t)$-th horizontal decomposition step induced by $b_{2t-1}=a_{2t}b_{2t}+b_{2t+1}$, the current residual rectangle of width $b_{2t-1}$ is partitioned, from left to right, into $a_{2t}$ squares of side length $b_{2t}$ and one residual sub-rectangle of width $b_{2t+1}$.
-Accordingly, we define the local column coordinate recursively by
-$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t}\quad(t\geq 1),$ (13)
-so that $j^{(2t)}$ is the remainder of $j^{(2t-2)}$ upon division by $b_{2t}$.
-Key observation: the column $a_{0}s+j$ enters the (2t1)(2t\!-\!1)-st vertical decomposition step if and only if, at each preceding horizontal decomposition step $(2k)$ with $k\in\{1,\dots,t-1\}$, the column lies in the residual sub-rectangle (rather than in one of the $a_{2k}$ squares of side length $b_{2k}$), i.e.,
+    which tiles the current residual rectangle of width $b_{2t-1}$ with $a_{2t}$ squares of side length $b_{2t}$ placed horizontally, leaving a residual rectangle of width $b_{2t+1}$.
+Fix $j\in\{0,1,\dots,b_{1}-1\}$ and track column $a_{0}s+j$ through this recursion.
+At stage $2t$ of the recursion, let $j^{(2t)}$ denote the column index of $a_{0}s+j$ within the current residual rectangle.
+Since vertical steps preserve the horizontal extent of the residual rectangle, $j^{(2t)}$ changes only at horizontal steps.
+Specifically, we define recursively
+$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}\pmod{b_{2t}}=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t},\quad t\geq 1.$ (14)
+The recursion proceeds deterministically.
+Since the Euclidean algorithm consists of exactly $m$ division steps, the trajectory of column $a_{0}s+j$ terminates at the first horizontal decomposition step $(2t)$ for which $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor<a_{2t}$, or exhausts the entire chain after at most $\lfloor m/2\rfloor$ horizontal decomposition steps.
+In either case, termination occurs after finitely many steps.
+Every intersection between column $a_{0}s+j$ and a Euclidean square in $\mathcal{P}(s,b_{1})$ occurs precisely at one of the following mutually exclusive events:
+  1. (i)
+    The column reaches a vertical decomposition step (2t1)(2t\!-\!1), i.e., it remains in the residual sub-rectangle after each preceding horizontal step $(2k)$ for $k=1,\dots,t-1$:
     $\forall k\in\{1,\dots,t-1\}:\ \Big\lfloor\frac{j^{(2k-2)}}{b_{2k}}\Big\rfloor=a_{2k}.$ (15)
-Then we have:
-  • (Vertical case: squares produced at the (2t1)(2t\!-\!1)-st vertical step).
-At the (2t1)(2t\!-\!1)-st vertical step, the recursion places $a_{2t-1}$ squares of side length $b_{2t-1}$ stacked vertically.
-Since these squares cover the full current width $b_{2t-1}$, the column intersects all of them, with the same within-square column index $v=j^{(2t-2)}$.
-Their upper-left row coordinates are
-    $y=\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\qquad q=0,1,\dots,a_{2t-1}-1.$
-    Hence $\mathcal{YV}(j)$ contains the $a_{2t-1}$ pairs
-    $(y,v)=\Bigg(\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\ j^{(2t-2)}\Bigg),\qquad q=0,1,\dots,a_{2t-1}-1.$
-  • (Horizontal terminating case: the unique square produced at the $(2t)$-th horizontal step). Now consider the subsequent $(2t)$-th horizontal step. If
-    $\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor<a_{2t},$
-    then the column lies in one of the $a_{2t}$ squares of side length $b_{2t}$ produced in this step, hence intersects exactly one such square. Its upper-left row coordinate equals
-    $y=\sum_{k=1}^{t}a_{2k-1}b_{2k-1},$
-    and the within-square column index is $v=j^{(2t)}$. Thus $\mathcal{YV}(j)$ contains the additional (terminating) pair
+    At such a step, the column intersects all $a_{2t-1}$ vertically stacked squares (since they cover the full width $b_{2t-1}$ of the current residual rectangle), yielding $a_{2t-1}$ distinct pairs in $\mathcal{YV}(j)$.
+  2. (ii)
+    The column terminates at a horizontal decomposition step $(2t)$, i.e., condition (15) holds for $k=1,\dots,t-1$ but
+    $\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor<a_{2t}.$
+    In this case the column lies within one of the $a_{2t}$ complete squares (rather than the residual sub-rectangle) and intersects exactly one square at this step, yielding a single additional pair in $\mathcal{YV}(j)$.
+No other intersections occur: when $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor=a_{2t}$ at a horizontal step, the column lies in the residual sub-rectangle of width $b_{2t+1}$ and does not intersect any square produced at that step; such intersections are accounted for in subsequent recursion steps.
+Consequently, $\mathcal{YV}(j)$ is precisely the disjoint union of contributions from events (i) over all $t$ satisfying (15) and from event (ii) at the unique terminating horizontal step (if any).
+Under condition (15), the contributions to $\mathcal{YV}(j)$ are as follows:
+  • Vertical contributions (step 2t12t\!-\!1). The column intersects all $a_{2t-1}$ vertically stacked squares of side length $b_{2t-1}$. Each such square has top-left row coordinate
+    $y=\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\quad q=0,1,\dots,a_{2t-1}-1,$
+    and local column index $v=j^{(2t-2)}$ within the square. Hence $\mathcal{YV}(j)$ contains the pairs
+    $(y,v)=\Bigg(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}+q\,b_{2t-1},\ j^{(2t-2)}\Bigg),\quad q=0,\dots,a_{2t-1}-1.$
+  • Horizontal terminating contribution (step $2t$).
+If condition (15) holds for $k=1,\dots,t-1$ but $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor<a_{2t}$, then the column lies within one of the $a_{2t}$ squares of side length $b_{2t}$ and contributes exactly one pair with top-left row coordinate
+    $y=\sum_{k=1}^{t}a_{2k-1}b_{2k-1}$
+    and local column index $v=j^{(2t)}$ within that square, namely
     $(y,v)=\Bigg(\sum_{k=1}^{t}a_{2k-1}b_{2k-1},\ j^{(2t)}\Bigg).$
-    If instead $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor=a_{2t}$, then the column lies in the residual sub-rectangle and this horizontal step contributes no pair, as the recursion continues to $(2t+1)$-th vertical decomposition.
-Collecting the contributions from the vertical case (for all reachable depths $t$) and from the horizontal terminating case (whenever termination occurs at step $(2t)$) yields the following closed form:
-$\mathcal{R}_{j}$ $=\alpha^{a_{0}s+j}\Bigg(\sum_{\begin{subarray}{c}1\leq t\leq\lceil m/2\rceil\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\end{subarray}}\ \sum_{q=0}^{a_{2t-1}-1}\beta^{\left(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\right)+q\,b_{2t-1}+j^{(2t-2)}}$
-$\hskip 73.97733pt+\,\sum_{\begin{subarray}{c}1\leq t\leq\lfloor m/2\rfloor\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\\ \left\lfloor\frac{j^{(2t-2)}}{b_{2t}}\right\rfloor<a_{2t}\end{subarray}}\beta^{\left(\sum_{k=1}^{t}a_{2k-1}b_{2k-1}\right)+j^{(2t)}}\Bigg),$
-which finished the proof of Proposition 1.
+    If instead $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor=a_{2t}$, the column remains in the residual rectangle; no square is intersected at this horizontal step, and the recursion proceeds to step $(2t+1)$.
+The closed-form expression for $\mathcal{R}_{j}$ aggregates precisely these two types of contributions:
+  • The first summation collects contributions from all vertical decomposition steps (2t1)(2t\!-\!1) that the column reaches (i.e., all $t$ satisfying (15)).
+At each such step, the column intersects all $a_{2t-1}$ squares produced, yielding $a_{2t-1}$ terms.
+  • The second summation captures the contribution from the unique terminating horizontal step $(2t)$ (if any), where the column falls into one of the $a_{2t}$ complete squares rather than the residual sub-rectangle, yielding exactly one additional term.
+No explicit partition of the range of $j$ is required: the constraints in the summation indices automatically select the relevant steps for each $j$ according to its trajectory through the Euclidean recursion.
+Aggregating all vertical contributions for indices $t$ satisfying (15) and the unique horizontal terminating contribution (if any) yields the closed-form expression:
+$\mathcal{R}_{j}$ $=\alpha^{a_{0}s+j}\Bigg(\sum_{\begin{subarray}{c}1\leq t\leq\lceil m/2\rceil\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor j^{(2k-2)}/b_{2k}\right\rfloor=a_{2k}\end{subarray}}\ \sum_{q=0}^{a_{2t-1}-1}\beta^{\left(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\right)+q\,b_{2t-1}+j^{(2t-2)}}$
+$\hskip 73.97733pt+\,\sum_{\begin{subarray}{c}1\leq t\leq\lfloor m/2\rfloor\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor j^{(2k-2)}/b_{2k}\right\rfloor=a_{2k}\\ \left\lfloor j^{(2t-2)}/b_{2t}\right\rfloor<a_{2t}\end{subarray}}\beta^{\left(\sum_{k=1}^{t}a_{2k-1}b_{2k-1}\right)+j^{(2t)}}\Bigg),$ (16)
+which completes the proof of Proposition 1.
 Appendix B Proof of Lemma 3
-Let
-$K=\sum_{u=0}^{s-1}\alpha^{u}S=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(u,v):(u,v)\in\mathcal{I}(s,p)\}.$
-From previous analysis, we only need to prove that for $(i,j)\in\mathcal{I}(s,b_{1})$, $R(i,a_{0}s+j)\in K$.
-Fix $(i,j)\in\mathcal{I}(s,b_{1})$, let $\tau(\cdot,\cdot)$ be as in Lemma 4, then
-$R(i,j):=\alpha^{a_{0}s+i+j}\beta^{\tau(i,a_{0}s+j)}.$
-Equivalently, if $(i,j)$ lies in the unique Euclidean square whose upper-left corner is $(y,x)$, then Lemma 4 gives $\tau(i,a_{0}s+j)=y+(j-x)=j+(y-x)$, hence
-$R(i,j)=\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{\,y-x}.$
-Since $\mathcal{R}_{j}\in S$, for every $i\in[0,s-1]$ we have $\alpha^{i}\mathcal{R}_{j}\in K$. By Proposition 1 together with Lemma  4, the element $\alpha^{i}\mathcal{R}_{j}$ has the form
-$\alpha^{i}\mathcal{R}_{j}=\alpha^{a_{0}s+i}(\alpha\beta)^{j}\Big(\beta^{y-x}+\sum_{\begin{subarray}{c}Q\text{ intersects col }j\\ Q\neq Q_{0}\end{subarray}}\beta^{y_{Q}-x_{Q}}\Big),$
-where $Q_{0}$ is the Euclidean square containing $(i,j)$ and, for each other intersected square $Q$, $(y_{Q},x_{Q})$ denotes its upper-left corner. Consequently,
-$R(i,j)=\alpha^{i}\mathcal{R}_{j}-\alpha^{a_{0}s+i}(\alpha\beta)^{j}\sum_{\begin{subarray}{c}Q\text{ intersects col }j\\ Q\neq Q_{0}\end{subarray}}\beta^{y_{Q}-x_{Q}}.$ (15)
-Therefore it is enough to show that for every $Q\neq Q_{0}$ intersecting column $j$,
-$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}\in K.$ (16)
-We prove (16) by induction on the column index $j$ ordered by the sequence of quotients appearing in the Euclidean recursion.
-Let $j^{(0)}:=j$ and $j^{(2t)}$ be defined by (13).
-Order columns by the lexicographic order of the finite sequence
-$\Big(\Big\lfloor\frac{j^{(0)}}{b_{2}}\Big\rfloor,\ \Big\lfloor\frac{j^{(2)}}{b_{4}}\Big\rfloor,\ \Big\lfloor\frac{j^{(4)}}{b_{6}}\Big\rfloor,\ \ldots\Big),$
-where the sequence stops when the reachability condition 14 first fails.
-Assume that for every column $j^{\prime}$ preceding $j$ in this order and every $i^{\prime}\in[0,s-1]$, we already have $R(i^{\prime},j^{\prime})\in K$.
-Fix an intersected square $Q\neq Q_{0}$ in column $j$.
-If $Q$ is a terminating square arising from a horizontal decomposition step, then its upper-left column coordinate is of the form $x_{Q}\equiv 0\pmod{b_{2t}}$ for the corresponding step size $b_{2t}$, and the contribution of $Q$ in $\alpha^{i}\mathcal{R}_{j}$ is the term $\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}$ (this is exactly the terminating term in (1)).
-Choose an integer $z\geq 1$ such that $(i+zb_{2t},\,j-zb_{2t})\in\mathcal{I}(s,b_{1})$ lies in the same terminating square pattern at that step (such a $z$ exists because the terminating square has side length $b_{2t}$ and column $j$ lies in a block with $\lfloor j^{(2t-2)}/b_{2t}\rfloor>0$ at that step).
-Then, by the definition of $r(\cdot,\cdot)$ via Lemma 4 and the identity
-$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=\alpha^{a_{0}s+i+zb_{2t}}(\alpha\beta)^{j-zb_{2t}}\beta^{y_{Q}-(x_{Q}-zb_{2t})}=\alpha^{a_{0}s+(i+zb_{2t})+(j-zb_{2t})}\beta^{\tau(i+zb_{2t},a_{0}s+j-zb_{2t})},$
-we obtain
-$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=R(i+zb_{2t},\,j-zb_{2t}).$
-Moreover, $j-zb_{2t}$ precedes $j$ in the above lexicographic order because the quotient $\lfloor j^{(2t-2)}/b_{2t}\rfloor$ strictly decreases while all earlier quotients remain unchanged.
-Hence, by the induction hypothesis, $R(i+zb_{2t},\,j-zb_{2t})\in K$, which proves (16) for terminating squares.
-If $Q$ is a vertical square in the Euclidean partition, then its contribution in (1) belongs to one of the vertical progressions at some vertical step, and the corresponding exponent difference $y_{Q}-x_{Q}$ differs from the target difference yxy-x by a positive multiple of the vertical step size at that stage.
-In this case, the element $\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}$ can be rewritten as
-$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=(\alpha\beta)^{j+(y_{Q}-x_{Q})}\alpha^{a_{0}s+i-(y_{Q}-x_{Q})}.$
-Since $Q\neq Q_{0}$ and $(i,j)\in Q_{0}$, the row index $i$ lies strictly above the vertical block indexed by $Q$ at that step, hence $0\leq a_{0}s+i-(y_{Q}-x_{Q})\leq a_{0}s-1$.
-Therefore $(\alpha\beta)^{j+(y_{Q}-x_{Q})}\alpha^{a_{0}s+i-(y_{Q}-x_{Q})}\in K$ by $\sum_{i=0}^{s-1}\alpha^{i}S^{(1)}\in K$, proving (16) for vertical squares.
-Thus every term in the sum on the right-hand side of (15) belongs to $K$, so (15) implies $R(i,j)\in K$.
-By induction over the above column order, we have $R(i,j)\in K$ for all $(i,j)\in\mathcal{I}(s,b_{1})$, which finished the proof.
+Setup and goal. Recall the Euclidean chain
+$p=a_{0}s+b_{1},\quad s=a_{1}b_{1}+b_{2},\quad b_{1}=a_{2}b_{2}+b_{3},\ \dots,\ b_{m-2}=a_{m-1}b_{m-1}+b_{m},\quad b_{m-1}=a_{m}b_{m},$ (17)
+where $b_{m}=\gcd(p,s)$.
+Let $B$ be the $p\times s$ array with $B(i,j)=\alpha^{i}(\alpha\beta)^{j}$ for $(i,j)\in[0,p-1]\times[0,s-1]$.
+Let $R=\mathsf{Reshape}_{p,s}(B)$ and $\overline{R}=\mathrm{Int}(R)$.
+By Lemma 4, for every $(i,j)\in[0,s-1]\times[0,p-1]$,
+$R(i,j)=\alpha^{i+j}\beta^{\tau(i,j)}$ (18)
+for some $\tau(i,j)\in[0,s-1]$, and for each fixed column $j$ the $\alpha$-exponents form the consecutive set $\{j,j+1,\dots,j+s-1\}$.
+Define
+$S:=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(0,t):0\leq t\leq p-1\},\qquad K:=\sum_{u=0}^{s-1}\alpha^{u}S.$
+Since $\mathrm{Int}$ acts as identity on the first $a_{0}s$ columns, $R(i,j)=\overline{R}(i,j)\in K$ for all $(i,j)\in[0,s-1]\times[0,a_{0}s-1]$.
+It remains to prove that for every $(i,j)\in[0,s-1]\times[0,b_{1}-1]$,
+$r(i,j):=R(i,a_{0}s+j)\in K.$ (19)
+B-A Proof roadmap: reduction to interference monomials
+For $0\leq j\leq b_{1}-1$, let $\mathcal{R}_{j}:=\overline{R}(0,a_{0}s+j)$ and let $\mathsf{Exp}(j)$ be the set of $\beta$-exponents appearing in column $a_{0}s+j$ of $R$.
+By definition of $\mathrm{Int}$, for every $0\leq i\leq s-1$,
+$\overline{R}(i,a_{0}s+j)=\alpha^{i}\mathcal{R}_{j}=\alpha^{a_{0}s+i+j}\sum_{t\in\mathsf{Exp}(j)}\beta^{t}\in K.$ (20)
+Since
+$r(i,j)=\alpha^{a_{0}s+i+j}\beta^{\tau(i,a_{0}s+j)},$
+we have
+$r(i,j)=\overline{R}(i,a_{0}s+j)-\alpha^{a_{0}s+i+j}\sum_{t\in\mathsf{Exp}(j)\setminus\{\tau(i,a_{0}s+j)\}}\beta^{t}.$ (21)
+Hence it is enough to prove
+$\alpha^{a_{0}s+i+j}\beta^{t}\in K\quad\forall t\in\mathsf{Exp}(j)\setminus\{\tau(i,a_{0}s+j)\}.$
+We will prove this by induction on a well-founded order $\prec$ and the split:
+  1. (C1)
+    $t<t^{\star}$;
+  2. (C2)
+    $t>t^{\star}$ and $i+j<t$;
+  3. (C3)
+    $t>t^{\star}$ and $i+j\geq t$.
+where $t^{\star}:=\tau(i,a_{0}s+j)$.
+B-B Quotient sequences and induction order
+Let $r:=\lfloor m/2\rfloor$ and $r^{\prime}:=\lceil m/2\rceil$. The last $b_{1}$ columns of $R$, denoted $R_{1}$, inherit the Euclidean square partition $\mathcal{P}(s,b_{1})$.
+For $j\in[0,b_{1}-1]$, define the remainder sequence
+$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}\pmod{b_{2t}}\quad(1\leq t\leq r),$
+and the even-step quotients $y_{2t}(j):=\big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\big\rfloor\in[0,a_{2t}]$. The column quotient sequence is
+$\mathbf{Y}(j):=\big(y_{2}(j),y_{4}(j),\dots,y_{2r}(j)\big).$ (22)
+For $i\in[0,s-1]$, define the remainder sequence
+$i^{(0)}:=i,\qquad i^{(2t)}:=i^{(2t-2)}\pmod{b_{2t-1}}\quad(1\leq t\leq r^{\prime}),$
+and the odd-step quotients $x_{2t-1}(i):=\Big\lfloor\frac{i^{(2t-2)}}{b_{2t-1}}\Big\rfloor\in[0,a_{2t-1}].$ The row quotient sequence is
+$\mathbf{X}(i):=\big(x_{1}(i),x_{3}(i),\dots,x_{2r^{\prime}-1}(i)\big).$ (23)
+Define a total order $\prec$ on $0,s-1]\times[0,b_{1}-1]$ by
+$(i^{\prime},j^{\prime})\prec(i,j)\quad\Longleftrightarrow\quad\big(\mathbf{Y}(j^{\prime}),\mathbf{X}(i^{\prime}),j^{\prime},i^{\prime}\big)<_{\rm lex}\big(\mathbf{Y}(j),\mathbf{X}(i),j,i\big).$
+In particular, we will use the column-priority consequence
+$j<j^{\prime}\Longrightarrow(i,j)\prec(i^{\prime},j^{\prime}).$
+This is a well-founded total order on the finite set $0,s-1]\times[0,b_{1}-1]$.
+B-C Two algebraic facts on K
+Fact 1 (Initial segment containment).
+For all $0\leq u\leq a_{0}s-1$ and $0\leq\theta\leq s-1$, $\alpha^{u}(\alpha\beta)^{\theta}\in K$.
+Proof.
+The elements $\overline{R}(0,qs+\theta)=\alpha^{qs}(\alpha\beta)^{\theta}\in S\subseteq K$ for $0\leq q\leq a_{0}-1$ and $0\leq\theta\leq s-1$.
+Multiplying by $\alpha^{v}$ for $0\leq v\leq s-1$ and using $K=\sum_{v=0}^{s-1}\alpha^{v}S$ yields $\alpha^{qs+v}(\alpha\beta)^{\theta}\in K$ for all $0\leq qs+v\leq a_{0}s-1$.
+∎
+Fact 2 (Field generation from consecutive powers).
+Let $\theta\in[0,s-1]$.
+If there exists an integer $z$ such that $\alpha^{z}\beta^{\theta},\alpha^{z+1}\beta^{\theta},\dots,\alpha^{z+p-1}\beta^{\theta}\in K$, then $\mathbb{E}\beta^{\theta}\subseteq K$.
+Proof.
+Since $\mathbb{E}:\mathbb{F}]=p$ and $\alpha$ generates $\mathbb{E}$ over $\mathbb{F}$, $\{\alpha^{z},\dots,\alpha^{z+p-1}\}$ is an $\mathbb{F}$-basis of $\mathbb{E}$. Hence
+$\operatorname{Span}_{\mathbb{F}}\{\alpha^{z}\beta^{\theta},\dots,\alpha^{z+p-1}\beta^{\theta}\}=\mathbb{E}\beta^{\theta}\subseteq K.$
+B-D Classification of Exp(j)
+From Proposition 1, each $t\in\mathsf{Exp}(j)$ has exactly one of the following forms.
+  1. 1.
+    Vertical contribution with index $(t_{0},q)$. There exist $t_{0}\geq 1$ and $q\in[0,a_{2t_{0}-1}-1]$ such that $y_{2k}(j)=a_{2k}$ for all $k\in\{1,\dots,t_{0}-1\}$ and
+    $t=\underbrace{\sum_{k=1}^{t_{0}-1}a_{2k-1}b_{2k-1}}_{=:Y_{t_{0}}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)}.$ (24)
+  2. 2.
+    Horizontal terminating contribution with index $t_{0}$. There exists $t_{0}\geq 1$ with $y_{2k}(j)=a_{2k}$ for all $k\in\{1,\dots,t_{0}-1\}$ and $y_{2t_{0}}(j)<a_{2t_{0}}$, such that
+    $t=\underbrace{\sum_{k=1}^{t_{0}}a_{2k-1}b_{2k-1}}_{=:L_{t_{0}}}+j^{(2t_{0})}.$ (25)
+Lemma 5 (Horizontal exponent is maximal in $\mathsf{Exp}(j)$).
+For every $(i,j)\in\mathcal{I}(s,b_{1})$, let $t^{\star}:=\tau(i,a_{0}s+j)$. Then
+[…]
+Proof.
+This is a direct corollary of the closed form of $\mathsf{Exp}(j)$ in Proposition 1.
+B-E Definition of a locating function
+Fix $\theta\in[0,s-1]$ and $k\in[0,b_{1}-1]$. Let
+$Q_{\theta,k}=[y_{\theta,k},\,y_{\theta,k}+\lambda_{\theta,k}-1]\times[x_{\theta,k},\,x_{\theta,k}+\lambda_{\theta,k}-1]$
+be the unique square in $\mathcal{P}(s,b_{1})$ containing $(\theta,k)$. Define
+$\Phi_{\theta}(k):=(i_{\theta,k},j_{\theta,k}):=\bigl(y_{\theta,k}+(k-x_{\theta,k}),\;x_{\theta,k}+(\theta-y_{\theta,k})\bigr).$ (26)
+Lemma 6 (Predecessors locating function $\Phi_{\theta}$).
+For every $\theta\in[0,s-1]$ and $k\in[0,b_{1}-1]$,
+$\tau(i_{\theta,k},a_{0}s+j_{\theta,k})=\theta,\qquad r(i_{\theta,k},j_{\theta,k})=\alpha^{a_{0}s+\theta+k}\beta^{\theta}.$
+Equivalently,
+$r\bigl(\Phi_{\theta}(k)\bigr)=\alpha^{a_{0}s+\theta+k}\beta^{\theta}.$
+Proof.
+Because $(\theta,k)\in Q_{\theta,k}$, there exist unique integers
+$a:=\theta-y_{\theta,k},\qquad b:=k-x_{\theta,k},$
+with $0\leq a,b\leq\lambda_{\theta,k}-1$. Hence $\Phi_{\theta}(k)=(y_{\theta,k}+b,\;x_{\theta,k}+a)\in Q_{\theta,k}$. Let
+$Q=[y,y+\lambda-1]\times[x,x+\lambda-1]\in\mathcal{P}(s,b_{1}).$
+In global column coordinates, the corresponding square of $R$ is
+$Q^{\prime}=[y,y+\lambda-1]\times[a_{0}s+x,\;a_{0}s+x+\lambda-1].$
+For $(u,v)\in Q$, write $u=y+r$, $v=x+c$ with $0\leq r,c\leq\lambda-1$. Then $(u,a_{0}s+v)\in Q^{\prime}$, so by Lemma 4 (applied on $Q^{\prime}$),
+$\tau(u,a_{0}s+v)=y+c=y+(v-x).$
+Applying this with $Q=Q_{\theta,k}$ and $(u,v)=\Phi_{\theta}(k)$ gives
+$\tau(i_{\theta,k},a_{0}s+j_{\theta,k})=y_{\theta,k}+\big(j_{\theta,k}-x_{\theta,k}\big)=y_{\theta,k}+a=\theta.$
+Therefore
+$r(i_{\theta,k},j_{\theta,k})=\alpha^{a_{0}s+i_{\theta,k}+j_{\theta,k}}\beta^{\theta}.$
+Finally,
+$i_{\theta,k}+j_{\theta,k}=\big(y_{\theta,k}+b\big)+\big(x_{\theta,k}+a\big)=(a+b)+(x_{\theta,k}+y_{\theta,k})=\theta+k,$
+so $r(i_{\theta,k},j_{\theta,k})=\alpha^{a_{0}s+\theta+k}\beta^{\theta}$. ∎
+B-F Lemmas for locating predecessors (C1/C3)
+Lemma 7 (Locating predecessors for Case C1).
+Fix $(i,j)\in[0,s-1]\times[0,b_{1}-1]$ and let
+$t\in\mathsf{Exp}(j)\setminus\{\tau(i,a_{0}s+j)\},\qquad t<t^{\star}:=\tau(i,a_{0}s+j).$
+Then $t$ is vertical (by Lemma 5). Equivalently, there exist $t_{0}\geq 1$ and $q\in[0,a_{2t_{0}-1}-1]$ such that
+$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)}.$
+Moreover, $q<x_{2t_{0}-1}(i)$. Then, for every $k\in[0,b_{1}-1]$, the point $\Phi_{t}(k)=(i_{k},j_{k})$ satisfies
+$(i_{k},j_{k})\prec(i,j).$
+Proof.
+By Lemma 5, $t$ is vertical, hence
+$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)}.$
+We now prove $q<x_{2t_{0}-1}(i)$.
+If $x_{2t_{0}-1}(i)\leq a_{2t_{0}-1}-1$, then
+$t^{\star}=\tau(i,a_{0}s+j)=Y_{t_{0}}+x_{2t_{0}-1}(i)\,b_{2t_{0}-1}+j^{(2t_{0}-2)}.$
+Comparing with
+$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)}$
+and using $t<t^{\star}$, we get $q<x_{2t_{0}-1}(i)$.
+If $x_{2t_{0}-1}(i)=a_{2t_{0}-1}$, then
+$q\leq a_{2t_{0}-1}-1<x_{2t_{0}-1}(i),$
+so again $q<x_{2t_{0}-1}(i)$.
+Therefore
+$q<x_{2t_{0}-1}(i).$ (27)
+Set
+$C:=\sum_{h=1}^{t_{0}-1}a_{2h}b_{2h}=b_{1}-b_{2t_{0}-1}.$
+From $y_{2h}(j)=a_{2h}$ for $h<t_{0}$, we have
+$j=C+j^{(2t_{0}-2)}\geq C.$
+Case I: $0\leq k\leq C-1$. Write $\Phi_{t}(k)=(i_{k},j_{k})$ and let
+$Q_{t,k}=[y_{t,k},y_{t,k}+\lambda_{t,k}-1]\times[x_{t,k},x_{t,k}+\lambda_{t,k}-1]$
+be the unique square containing $(t,k)$.
+Since $k<C$ and $C=b_{1}-b_{2t_{0}-1}$, the square $Q_{t,k}$ is contained in $0,s-1]\times[0,C-1]$.
+Indeed, $x=C$ is a square-edge boundary after horizontal step $(2t_{0}-2)$; all later decomposition steps act only inside the residual strip $C,b_{1}-1]$.
+Since $k<C$, the point $(t,k)$ lies strictly to the left of this boundary, so its containing square cannot cross $x=C$.
+Hence
+$x_{t,k}+\lambda_{t,k}\leq C.$
+By (26),
+$j_{k}=x_{t,k}+\bigl(t-y_{t,k}\bigr)\leq x_{t,k}+\lambda_{t,k}-1\leq C-1<j,$
+where the first inequality uses $(t,k)\in Q_{t,k}$, which implies $0\leq t-y_{t,k}\leq\lambda_{t,k}-1$. Hence $j_{k}<j$, so by the column-priority consequence of $\prec$,
+$(i_{k},j_{k})\prec(i,j).$
+Case II: $C\leq k\leq b_{1}-1$. From
+$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)},\qquad 0\leq j^{(2t_{0}-2)}\leq b_{2t_{0}-1}-1,$
+we get
+$Y_{t_{0}}+q\,b_{2t_{0}-1}\leq t\leq Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1.$
+Also, $C\leq k\leq b_{1}-1$ and $C=b_{1}-b_{2t_{0}-1}$ imply
+$C\leq k\leq C+b_{2t_{0}-1}-1.$
+Hence
+$(t,k)\in Q^{\ast},$
+where
+$Q^{\ast}=[Y_{t_{0}}+q\,b_{2t_{0}-1},\,Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1]\times[C,\,b_{1}-1].$
+Hence $x_{t,k}=C$, $y_{t,k}=Y_{t_{0}}+q\,b_{2t_{0}-1}$, $\lambda_{t,k}=b_{2t_{0}-1}$ and
+$t-y_{t,k}=j^{(2t_{0}-2)}.$
+Therefore
+$j_{k}=x_{t,k}+\bigl(t-y_{t,k}\bigr)=C+j^{(2t_{0}-2)}=j.$
+Also
+$i_{k}=y_{t,k}+(k-C)\in[Y_{t_{0}}+q\,b_{2t_{0}-1},\,Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1],$
+which implies
+$x_{2t_{0}-1}(i_{k})=q.$
+Combining this with (27), we obtain
+$x_{2t_{0}-1}(i_{k})<x_{2t_{0}-1}(i).$
+For odd indices $1,3,\dots,2t_{0}-3$, the quotients of $i_{k}$ and $i$ coincide; hence
+$\mathbf{X}(i_{k})<_{\rm lex}\mathbf{X}(i).$
+Since $j_{k}=j$, this gives $(i_{k},j_{k})\prec(i,j)$.
+Combining Case I and Case II yields $(i_{k},j_{k})\prec(i,j)$ for all $k\in[0,b_{1}-1]$. ∎
+Lemma 8 (Locating a predecessor for Case C3).
+Let $(i,j)\in[0,s-1]\times[0,b_{1}-1]$, let
+$t^{\star}:=\tau(i,a_{0}s+j),\qquad t\in\mathsf{Exp}(j)\setminus\{t^{\star}\},$
+and assume
+$t>t^{\star},\qquad i+j\geq t.$
+Define
+$\delta:=i+j-t.$
+Then:
+  1. (a)
+    $\delta\in[0,b_{1}-1]$.
+  2. (b)
+    The point $(i_{\delta},j_{\delta}):=\Phi_{t}(\delta)$ satisfies
+    $r(i_{\delta},j_{\delta})=\alpha^{a_{0}s+i+j}\beta^{t},\qquad(i_{\delta},j_{\delta})\prec(i,j).$
+Proof.
+Proof of (a). Since $i+j\geq t$, we have $\delta\geq 0$. It remains to show $\delta\leq b_{1}-1$.
+Case A1: $t$ is vertical. Write
+$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)},\qquad C:=\sum_{h=1}^{t_{0}-1}a_{2h}b_{2h}=b_{1}-b_{2t_{0}-1},\qquad j=C+j^{(2t_{0}-2)}.$
+Let $Q_{t,j}$ be the unique square in $\mathcal{P}(s,b_{1})$ containing $(t,j)$. From $0\leq j^{(2t_{0}-2)}\leq b_{2t_{0}-1}-1$, we have
+$Y_{t_{0}}+q\,b_{2t_{0}-1}\leq t\leq Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1,\qquad C\leq j\leq b_{1}-1.$
+Hence $(t,j)$ lies in
+$Q_{v}:=[Y_{t_{0}}+q\,b_{2t_{0}-1},\,Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1]\times[C,b_{1}-1].$
+At vertical step $(2t_{0}-1)$, $Q_{v}$ is one square of $\mathcal{P}(s,b_{1})$; therefore $Q_{t,j}=Q_{v}$. By Lemma 4, for all
+$u\in[Y_{t_{0}}+q\,b_{2t_{0}-1},\,Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1],$
+we have
+$\tau(u,a_{0}s+j)=Y_{t_{0}}+q\,b_{2t_{0}-1}+(j-C)=t.$
+For fixed $j$, Proposition 1 orders exponents by Euclidean terms, so rows below this interval satisfy $\tau(u,a_{0}s+j)>t$.
+Since $t^{\star}=\tau(i,a_{0}s+j)<t$, row $i$ cannot be in this interval or below it.
+Hence
+$i\leq Y_{t_{0}}+q\,b_{2t_{0}-1}-1.$
+Therefore
+$\delta=i+j-t=i+C-\bigl(Y_{t_{0}}+q\,b_{2t_{0}-1}\bigr)\leq C-1\leq b_{1}-1.$
+Case A2: $t$ is horizontal. Write
+$t=L_{t_{0}}+j^{(2t_{0})},$
+where
+$C_{h}:=\sum_{h=1}^{t_{0}-1}a_{2h}b_{2h}+y_{2t_{0}}(j)b_{2t_{0}}.$
+By the remainder decomposition,
+$j=\sum_{h=1}^{t_{0}-1}a_{2h}b_{2h}+j^{(2t_{0}-2)},\qquad j^{(2t_{0}-2)}=y_{2t_{0}}(j)b_{2t_{0}}+j^{(2t_{0})}.$
+Hence
+$j=C_{h}+j^{(2t_{0})}.$
+Let $Q_{t,j}$ be the unique square in $\mathcal{P}(s,b_{1})$ containing $(t,j)$. From $0\leq j^{(2t_{0})}\leq b_{2t_{0}}-1$, we have $(t,j)\in Q_{h}$, where
+$Q_{h}:=[L_{t_{0}},\,L_{t_{0}}+b_{2t_{0}}-1]\times[C_{h},\,C_{h}+b_{2t_{0}}-1].$
+At horizontal step $(2t_{0})$, $Q_{h}$ is one square of $\mathcal{P}(s,b_{1})$; hence $Q_{t,j}=Q_{h}$. By Lemma 4, for all $u\in[L_{t_{0}},L_{t_{0}}+b_{2t_{0}}-1]$,
+$\tau(u,a_{0}s+j)=L_{t_{0}}+(j-C_{h})=L_{t_{0}}+j^{(2t_{0})}=t.$
+By Euclidean recursion,
+$s=b_{0}=\sum_{h=1}^{t_{0}}a_{2h-1}b_{2h-1}+b_{2t_{0}}=L_{t_{0}}+b_{2t_{0}},$
+so $L_{t_{0}},L_{t_{0}}+b_{2t_{0}}-1]=[L_{t_{0}},s-1]$. If $i\geq L_{t_{0}}$, then $\tau(i,a_{0}s+j)=t$, contradicting $t^{\star}<t$. Hence $i\leq L_{t_{0}}-1$, and
+$\delta=i+j-t=i+C_{h}-L_{t_{0}}\leq C_{h}-1.$
+Since $j=C_{h}+j^{(2t_{0})}$ and $0\leq j\leq b_{1}-1$, we have $C_{h}\leq b_{1}-1$, therefore
+$\delta\leq b_{1}-1.$
+Combining A1 and A2, we get $\delta\in[0,b_{1}-1]$.
+Proof of (b). Set $(i_{\delta},j_{\delta}):=\Phi_{t}(\delta)$. By Lemma 6,
+$r(i_{\delta},j_{\delta})=\alpha^{a_{0}s+t+\delta}\beta^{t}=\alpha^{a_{0}s+i+j}\beta^{t}.$
+It remains to prove $(i_{\delta},j_{\delta})\prec(i,j)$.
+Case B1: $t$ is vertical. From part (a), $\delta\leq C-1$. Let
+$Q_{t,\delta}=[y_{t,\delta},y_{t,\delta}+\lambda_{t,\delta}-1]\times[x_{t,\delta},x_{t,\delta}+\lambda_{t,\delta}-1]$
+be the unique square containing $(t,\delta)$. Since $(t,\delta)\in Q_{t,\delta}$,
+$0\leq t-y_{t,\delta}\leq\lambda_{t,\delta}-1,\qquad 0\leq\delta-x_{t,\delta}\leq\lambda_{t,\delta}-1.$
+The line $x=C$ is a square-edge boundary after step $2t_{0}-2$. Because $\delta<C$, $Q_{t,\delta}$ cannot cross this boundary, hence
+$x_{t,\delta}+\lambda_{t,\delta}\leq C.$
+Therefore
+$j_{\delta}=x_{t,\delta}+(t-y_{t,\delta})\leq x_{t,\delta}+\lambda_{t,\delta}-1\leq C-1.$
+Also $j=C+j^{(2t_{0}-2)}$, so
+$j\geq C.$
+Therefore $j_{\delta}\leq C-1<j$, and by the column-priority consequence of $\prec$,
+$(i_{\delta},j_{\delta})\prec(i,j).$
+Case B2: $t$ is horizontal. From part (a), $\delta\leq C_{h}-1$. Let
+$Q_{t,\delta}=[y_{t,\delta},y_{t,\delta}+\lambda_{t,\delta}-1]\times[x_{t,\delta},x_{t,\delta}+\lambda_{t,\delta}-1]$
+be the unique square containing $(t,\delta)$. Since $(t,\delta)\in Q_{t,\delta}$,
+$0\leq t-y_{t,\delta}\leq\lambda_{t,\delta}-1,\qquad 0\leq\delta-x_{t,\delta}\leq\lambda_{t,\delta}-1.$
+The line $x=C_{h}$ is a square-edge boundary at horizontal step $2t_{0}$. Because $\delta<C_{h}$, $Q_{t,\delta}$ cannot cross this boundary, so
+$x_{t,\delta}+\lambda_{t,\delta}\leq C_{h}.$
+Hence
+$j_{\delta}=x_{t,\delta}+(t-y_{t,\delta})\leq C_{h}-1.$
+Write $j=C_{h}+j^{(2t_{0})}$. Since $j\geq C_{h}$, we have $j_{\delta}\leq C_{h}-1<j$, hence by the column-priority consequence of $\prec$,
+$(i_{\delta},j_{\delta})\prec(i,j).$
+B-G Main induction
+We prove (19) for all $(i,j)\in[0,s-1]\times[0,b_{1}-1]$ by induction on $\prec$.
+Initialization.
+For $j=0$ and $i\in[0,b_{1}-1]$, every interference exponent $t\in\mathsf{Exp}(0)\setminus\{0\}$ satisfies $t\geq b_{1}>i=i+j$. Hence
+$\alpha^{a_{0}s+i}\beta^{t}=(\alpha\beta)^{t}\alpha^{a_{0}s+i-t}\in K$
+by Fact 1, so $r(i,0)\in K$ on this initial boundary set.
+Inductive step.
+Fix $(i,j)$ and assume
+$r(i^{\prime},j^{\prime})\in K\quad\forall(i^{\prime},j^{\prime})\prec(i,j).$
+Set $t^{\star}:=\tau(i,a_{0}s+j)$, and take any $t\in\mathsf{Exp}(j)\setminus\{t^{\star}\}$.
+We show
+$\alpha^{a_{0}s+i+j}\beta^{t}\in K.$
+Case C1: $t<t^{\star}$. By Lemma 5, this branch contains no horizontal term; thus $t$ is vertical and Lemma 7 applies.
+First, Fact 1 gives
+$\alpha^{t+u}\beta^{t}\in K,\qquad u=0,1,\dots,a_{0}s-1.$
+Next, for each $k\in[0,b_{1}-1]$, let $(i_{k},j_{k})=\Phi_{t}(k)$. By Lemma 7, $(i_{k},j_{k})\prec(i,j)$, so by induction hypothesis,
+$r(i_{k},j_{k})\in K.$
+By Lemma 6,
+$r(i_{k},j_{k})=\alpha^{a_{0}s+t+k}\beta^{t},$
+thus
+$\alpha^{a_{0}s+t+k}\beta^{t}\in K,\qquad k=0,1,\dots,b_{1}-1.$
+Therefore we have $p=a_{0}s+b_{1}$ consecutive monomials
+$\alpha^{t}\beta^{t},\alpha^{t+1}\beta^{t},\dots,\alpha^{t+p-1}\beta^{t}\in K.$
+By Fact 2,
+$\beta^{t}\mathbb{E}\subseteq K.$
+In particular, $\alpha^{a_{0}s+i+j}\beta^{t}\in K$.
+Case C2: $t>t^{\star}$ and $i+j<t$. Then
+$\alpha^{a_{0}s+i+j}\beta^{t}=(\alpha\beta)^{t}\alpha^{a_{0}s+i+j-t},$
+with $0\leq a_{0}s+i+j-t\leq a_{0}s-1$. Hence Fact 1 gives $\alpha^{a_{0}s+i+j}\beta^{t}\in K$.
+Case C3: $t>t^{\star}$ and $i+j\geq t$. Let $\delta:=i+j-t$. By part (a) of Lemma 8, $\delta\in[0,b_{1}-1]$. By part (b) of Lemma 8, with $(i_{\delta},j_{\delta})=\Phi_{t}(\delta)$,
+$r(i_{\delta},j_{\delta})=\alpha^{a_{0}s+i+j}\beta^{t},\qquad(i_{\delta},j_{\delta})\prec(i,j).$
+So induction gives $r(i_{\delta},j_{\delta})\in K$, i.e.
+$\alpha^{a_{0}s+i+j}\beta^{t}\in K.$
+Thus in Cases C1, C2 and C3 every interference monomial in (21) belongs to $K$. Together with (20), this yields $r(i,j)\in K$. The induction is complete.
+B-H Conclusion
+By induction over $\prec$, (19) holds for all $(i,j)\in[0,s-1]\times[0,b_{1}-1]$.
+Hence every entry of $R$ belongs to $K$.
+Since $R$ is a rearrangement of entries of $B$, we have $\mathbb{K}=\operatorname{Span}_{\mathbb{F}}(B)\subseteq K$.
+As $K\subseteq\mathbb{K}$ by definition, $K=\mathbb{K}$, namely
+$\mathbb{K}=\sum_{u=0}^{s-1}\alpha^{u}S.$
+$\square$
+B-I Concrete verification for $m=4$ (all $a_{i}=2$)
+Take p=70,s=29p=70,\ s=29, so the Euclidean chain is
+$70=2\cdot 29+12,\quad 29=2\cdot 12+5,\quad 12=2\cdot 5+2,\quad 5=2\cdot 2+1,\quad 2=2\cdot 1.$
+Hence
+$b_{1}=12,\ b_{2}=5,\ b_{3}=2,\ b_{4}=1,\qquad a_{0}s=58.$
+Exponent array on $R_{1}$ (last $b_{1}$ columns of $R$). Since $a_{0}s=58$, these columns are indexed by global columns $58+j$ with $0\leq j\leq 11$. For $(i,j)\in[0,28]\times[0,11]$, we have
+$\tau(i,58+j)=\begin{cases}j,&0\leq i\leq 11,\\ 12+j,&12\leq i\leq 23,\\ 24+(j\bmod 5),&24\leq i\leq 28,\ 0\leq j\leq 9,\\ 24+(j-10),&24\leq i\leq 25,\ j\in\{10,11\},\\ 26+(j-10),&26\leq i\leq 27,\ j\in\{10,11\},\\ 28,&i=28,\ j\in\{10,11\}.\end{cases}$
+For column $j=10$ in $R_{1}$ (global column index $58+10=68$), we have
+$j^{(0)}=10,\quad y_{2}(10)=\lfloor 10/5\rfloor=2,\quad j^{(2)}=0,\quad y_{4}(10)=0,\quad j^{(4)}=0.$
+Therefore
+$\mathsf{Exp}(10)=\{10,22,24,26,28\},$
+where $28$ is the unique horizontal term, and
+$\tau(i,68)=\begin{cases}10,&0\leq i\leq 11,\\ 22,&12\leq i\leq 23,\\ 24,&24\leq i\leq 25,\\ 26,&26\leq i\leq 27,\\ 28,&i=28.\end{cases}$
+Equivalently,
+$(\tau(0,68),\dots,\tau(28,68))=(\underbrace{10,\dots,10}_{12},\underbrace{22,\dots,22}_{12},24,24,26,26,28).$
+Five-segment repair on column $j=10$ of $R_{1}$. Let
+$I_{1}=[0,11],\quad I_{2}=[12,23],\quad I_{3}=[24,25],\quad I_{4}=[26,27],\quad I_{5}=\{28\}.$
+On these segments,
+$t^{\star}(i):=\tau(i,68)=10,22,24,26,28\quad\text{for }i\in I_{1},I_{2},I_{3},I_{4},I_{5}.$
+For each $i\in[0,28]$ and each $t\in\mathsf{Exp}(10)\setminus\{t^{\star}(i)\}$, we split by Cases C1, C2, C3.
+Segment $I_{1}$ ($t^{\star}=10$). For $t\in\{22,24,26,28\}$ and $i\in I_{1}$,
+$i+10\leq 21<t,$
+so this segment is entirely Case C2.
+Segment $I_{2}$ ($t^{\star}=22$). For $t=10$, this is Case C1. For $t\in\{24,26,28\}$, it is Case C2 when $i+10<t$, and Case C3 when $i+10\geq t$, i.e.
+[…]
+[…]
+[…]
+Segment $I_{3}$ ($t^{\star}=24$). $t\in\{10,22\}$ gives Case C1; $t\in\{26,28\}$ gives Case C3, since $i+10\geq 34$.
+Segment $I_{4}$ ($t^{\star}=26$). $t\in\{10,22,24\}$ gives Case C1; $t=28$ gives Case C3.
+Segment $I_{5}$ ($t^{\star}=28$). All $t\in\{10,22,24,26\}$ are Case C1.
+Thus every interference term in column $j=10$ is covered by one of C1/C2/C3. Case C2 is handled by Fact 1, Case C1 by Lemma 7, and Case C3 by Lemma 8.
+How one concrete entry is recovered: $(i,j)=(26,10)$. Here
+$t^{\star}=\tau(26,68)=26,\qquad\mathsf{Exp}(10)\setminus\{t^{\star}\}=\{10,22,24,28\}.$
+Since $26+10=36>28$, there is no C2 term; $t=10,22,24$ are C1 and $t=28$ is C3. We show explicitly that the C1 representative $t=24$ and the C3 term $t=28$ are both eliminable.
+C1 representative ($t=24$). For $k\in[0,11]$, let $(i_{k},j_{k})=\Phi_{24}(k)$. For example, take $k=8$. In this instance, $(24,8)$ lies in the square
+$Q_{24,8}=[24,28]\times[5,9],$
+so $y_{24,8}=24$, $x_{24,8}=5$, and by (26),
+$\Phi_{24}(8)=\big(24+(8-5),\;5+(24-24)\big)=(27,5)\prec(26,10).$
+In fact, by Lemma 7,
+$(i_{k},j_{k})\prec(26,10)\quad\text{for all }k.$
+By Lemma 6,
+$r(i_{k},j_{k})=\alpha^{58+24+k}\beta^{24},$
+so $\alpha^{58+24}\beta^{24},\dots,\alpha^{58+35}\beta^{24}\in K$. Together with the initial block from Fact 1, Fact 2 implies that $\beta^{24}\mathbb{E}\subseteq K$, hence
+$\alpha^{58+26+10}\beta^{24}\in K.$
+Therefore, the interference term with exponent $t=24$ at $(26,10)$ can be canceled.
+C3 term ($t=28$). Set $\delta=26+10-28=8$. By Lemma 8, part (b), with $(i_{\delta},j_{\delta})=\Phi_{28}(8)$,
+$r(i_{\delta},j_{\delta})=\alpha^{58+26+10}\beta^{28},\qquad(i_{\delta},j_{\delta})\prec(26,10).$
+To compute this point explicitly: $(28,8)$ is in
+$Q_{28,8}=[24,28]\times[5,9],$
+hence $y_{28,8}=24$, $x_{28,8}=5$, and by (26),
+$\Phi_{28}(8)=\big(24+(8-5),\;5+(28-24)\big)=(27,9).$
+Thus its column index is $9<10$, so this predecessor is already recovered. Hence
+$\alpha^{58+26+10}\beta^{28}\in K,$
+which cancels the $t=28$ interference at $(26,10)$.
+The remaining terms $t=10,22$ are also C1 and are handled identically to $t=24$. So every interference term at the specific entry $(26,10)$ is removable, exactly as required by the induction proof.